### v1

# Alternative Approaches for Computing Highest-Density Regions

Abstract
Many statistical problems require estimating a density function, say $f$, from data samples. In this work, for example, we are interested in highest-density regions (HDRs), i.e., minimum volume sets that contain a given probability. HDRs are typically computed using a density quantile approach, which, in the case of unknown densities, involves their estimation. This task turns out to be far from trivial, especially over increased dimensions and when data are sparse and exhibit complex structures (e.g., multimodalities or particular dependencies). We address this challenge by exploring alternative approaches to build HDRs that overcome direct (multivariate) density estimation. First, we generalize the density quantile method–currently implementable on the basis of a consistent estimator of the density–to neighbourhood measures, i.e., measures that preserve the order induced in the sample by $f$. Second, we discuss a number of suitable probabilistic- and distance-based measures such as the $k$-nearest neighbourhood Euclidean distance. Third, motivated by the ubiquitous role of copula modeling in modern statistics, we explore its use in the context of probabilistic-based measures. An extensive comparison among the introduced measures is provided, and their implications for computing HDRs in real-world problems are discussed.
Keywords Statistical intervals  $\cdot$ Density estimation  $\cdot$ Copula models  $\cdot$ k-nearest neighborhood
1 Introduction
A ubiquitous problem in statistics is to derive statistical intervals or regions–especially in the multivariate setting–for population parameters or other unknown quantities. Their role is to provide a way to quantify and describe the uncertainty about a quantity of interest, or simply a way to summarize the information contained in a distribution. Statistical regions may address different problems. For example, a confidence interval (CI) describes the uncertainty surrounding an estimate for an unknown parameter, while a prediction interval provides bounds for one or more future observations. Alternatively, a tolerance interval is the interval expected to contain a specified proportion of the sampled population. In a Bayesian setting, highest posterior density or credible regions provide, in a natural way, set estimates for a specific parameter (Box and Tiao,, 1992; Turkkan and Pham-Gia,, 1993). We refer to Meeker et al., (2017) and Krishnamoorthy and Mathew, (2009) for an overview. Furthermore, even when the focus is on a specific type of interval, e.g., a two-sided 95% prediction interval, questions on the way this should be defined may still arise. Should we use the interval symmetric about the mean, the interval of shortest length, or the interval that minimizes the probability of covering a given set? Although each of these intervals has 95% coverage, they may all be different. Consider, for example, the case of nonsymmetric or/and multimodal distributions, with an illustration given in Figure 1. An additional layer of complexity arises in multivariate settings, where there are no unique agreed definitions, and generalizations include the concept of simultaneous CIs (Guilbaud,, 2008), multivariate CIs (Korpela et al.,, 2017), or different definitions for multivariate quantiles (Figalli,, 2018; Cai,, 2010; Coblenz et al.,, 2018), among others.
The multivariate setting will be the target of this work, with a focus on bivariate distributions. In particular, our interest is devoted to statistical regions for summarizing probability distributions in the form of highest-density regions (HDRs; Hyndman,, 1996). Statistical regions other than HDRs are beyond the scope of the present work. As the name suggests, an HDR specifies the set of points of highest density: The density for points inside the region must be higher than that for points outside it. More specifically, considering a $d$-dimensional continuous variable of interest $X\in\mathbb{R}^{d}$ with probability density function $f$, the concrete problem is to estimate minimum volume sets of the form $C(f_{\alpha})=\{x\!:f(x)\geq f_{\alpha}\}$, such that $P(X\in C(f_{\alpha}))\geq 1-\alpha$, where $1-\alpha$, with $\alpha\in(0,1)$, represents a prespecified coverage probability. Although they share substantial similarities with multivariate quantiles, estimating an HDR differs from estimating level sets in that one is interested in specifying a probability content rather than the level directly. This complicates the problem, and we refer to Doss and Weng, (2018) for more details.
The scope of an HDR can be wide and diverse; the following are possible applications.
Forecasting
To obtain a prediction or forecast region for a set of observable variables in order to inform the most likely future realizations and convey in a simple way the accuracy of a forecast (for illustrative examples, see e.g., Hyndman,, 1996; Kim et al.,, 2011).
Anomaly detection
To detect abnormal observations from a sample: if a data point does not belong to a region of “normal” or concentrated data (the HDR), then it is regarded anomalous (see e.g., Steinwart et al.,, 2005, and references therein).
Unsupervised or semi-supervised classification
To identify areas or clusters with a relatively high concentration of a given phenomenon; see, e.g., the work of Saavedra-Nieves, (2022), aimed at finding areas of high incidence of coronavirus.
Due to their flexibility “to convey both multimodality and asymmetry”, HDRs are argued to be a more effective summary of the distribution (Hyndman,, 1995). In the case of unimodal symmetric distributions, such as the normal distribution, an HDR coincides with the usual probability region symmetric about the mean, spanning the $\alpha/2$ and $21-\alpha/$ 2 quantiles. However, in the case of a multimodal distribution, it may consist of several disjoint subregions, each containing a local mode. This provides useful information that is “masked” by other probability regions. An illustrative example is provided in Figure 1, where an HDR is compared to an equal-tailed interval.
Figure 1: Comparison between two $100(1-\alpha)\%$ probability intervals for a normal mixture density: an HDR and an equal-tailed interval. The coverage parameter $\alpha$ is set to $0.05$.
In principle, knowledge of an accurate estimator of the underlying probability density function $f$ of a variable $X$ is sufficient to compute its HDRs. However, existing methods for estimating $f$, such as the kernel density estimator (KDE; Parzen,, 1962) or the local likelihood approach (Hjort and Jones,, 1996), work very well for unidimensional problems, but may be inefficient for multidimensional problems (Liu et al.,, 2007). In fact, over increased dimensions, KDE suffers from the difficulty of finding optimal kernel functions and the corresponding bandwidths (i.e., the smoothing parameters). In particular, bandwidth selection in KDE is recognized as the most crucial and difficult step (see e.g., Chapter 2 in Wand and Jones, 1994a, ), with no definite and broadly accepted solution. As an illustrative example, in Figure 2 we estimate a $95\%$ HDR for a bivariate Gaussian distribution using KDE and three different bandwidth specifications. These are given in the works of Silverman, (1996)–in the form of rule of thumb, in Wand and Jones, 1994b –currently employed in the ks and hdrcde packages in R, and in Chacón et al., (2011)–under optimal asymptotic considerations. As depicted in Figure 2, the three different values result in three different HDRs, with an open question on the most appropriate one.
Figure 2: Estimated 95% HDR on a sample of $n=200$ units using KDE with three bandwidth specifications. Cadet-blue points are the HDR points, while purple x-marks and red triangles denote the points outside the HDR. The blue contour line represents the true level set (i.e., the one delimiting a 95% density region) of a bivariate Gaussian distribution (see Scenario S2 in Section 4).
In particular, if the task would be to identify abnormal or anomalous values, contradicting conclusions would arise from the three different HDRs (see e.g., the red triangle points in Figure 2). More concretely, our analysis is directly motivated by, and sees our involvement in, the World Anti-Doping Agency (WADA) mission of doping detection.
Furthermore, high-dimensional data also pose challenges from an algorithmic/computational perspective. Taken together, these aspects hamper the ability to derive an appropriate HDR. Motivated by these existing density-estimation issues, our research question is the following: “Can we relax the multivariate density estimation problem, either by using alternative ways (to direct density estimation, e.g., using distance-based proxies) that will robustly inform the density around a given point or by reducing the multidimensional estimation problem to a lower dimension?”. We address this question by exploring, evaluating, and proposing some novel approaches to building HDRs.
The contribution of the present work is fourfold. First, we generalize the density-quantile approach for computing HDRs–currently implementable on the basis of a consistent estimator of the density–to neighbourhood measures (Munoz and Moguerza,, 2006), i.e., measures that preserve the order induced in the sample by the density function (Section 2). Second, we elaborate on and evaluate a number of suitable probabilistic- and distance-based measures, including a variation of the measure adopted by the k-nearest neighbors algorithm (Section 3). Third, motivated by the ubiquitous role of copula modeling (Nelsen,, 2006) in modern statistics, among probabilistic-based measures, we explore the use of copulae in an HDR estimation context. This is done in Section 3.4. Specifically, copulae make it possible to relax the estimation of multivariate random vectors, by separately estimating the marginals and their dependence structure, i.e., the copula model. In addition, by placing a strong focus on the dependence model, a copula approach has the advantage of better capturing data specificities, especially when these exhibit complex relationships such as asymmetric and/or tail dependencies. Finally, in Section 4, we provide a comprehensive comparison between the introduced measures and standard density estimators in all simulated scenarios, as well as in the MAGIC data set, which classifies high-energy Gamma particles in the atmosphere (Bock et al.,, 2004). Empirical studies are focused on bivariate scenarios that vary according to the complexity of the data (marginal and dependence structure, multimodality, etc) and the sample size. A distribution on a compact support, namely, the Dirichlet distribution, is also considered. We conclude by summarizing the main findings and discussing their implications for computing HDRs in real-world problems.
2 HDR: Formulation and Computation
Let $\{X_{1},\dots,X_{n}\}$ be a sample of $n$ independent and identically distributed (iid) replications of a random variable $X$ defined on $\mathbb{R}^{d}$, with $d\geq 1$. In our problem, we assume to have access to a sample $s_{n}=\{x_{1},\dots,x_{n}\}\in S_{n}$ of their actual realizations, which we wish to use to build an HDR, that is, a statistical region containing those sample values of relatively high density (see Definition 1). We denote by $X_{ij}\ (x_{ij})$ the $j$-th component of $X_{i}\ (x_{i})$, $j=1,\dots,d$ and $i=1,\dots,n$. We restrict our discussion to continuous random variables $X$ and, unless otherwise stated, we denote by $f$ their probability density function (PDF) and by $F$ their cumulative density function (CDF).
Definition 1 (Highest-density region; Hyndman, (1996))
Denote by $f$ the PDF of a continuous, possibly multivariate, random variable $X\in\mathbb{R}^{d}$; the $100(1-\alpha)\%$ HDR is defined as the subset $C(f_{\alpha})$ of the sample space of $X$ such that:
$C(f_{\alpha})=\{x:f(x)\geq f_{\alpha}\},$
where $f_{\alpha}$ is the largest constant such that $P(X\in C(f_{\alpha}))\geq 1-\alpha$, with $\alpha\in(0,1)$.
One of the most distinctive properties of HDRs is that, among all regions of probability coverage $100(1-\alpha)\%$, the HDR has the smallest possible volume. The notion of “smallest” is to be understood with respect to some measure such as the usual Lebesgue measure; in the continuous one-dimensional case that would lead to the shortest-length set, while in two dimensions that would be the smallest-area set. It also follows from the definition that the boundary of an HDR consists of those values of the sample space with equal density. Hence a plot of a bivariate HDR has as the boundary a contour plot (see Figure 2).
2.1 Density quantile approach
The study of HDRs has been largely enhanced by Hyndman, who proposed the density quantile approach (outlined in Proposition 1) to compute multivariate HDRs (Hyndman,, 1996). Today, this still represents the typical strategy and involves estimating the density.
Proposition 1 (Hyndman, (1996))
Let $\{f(x_{1}),\dots,f(x_{m})\}$ be a sample of independent observations of size $m$ of the random variable $Y=f(X)$, with $f$ a bounded and continuous function in $x$. Consider the ordered sample $\{f_{(1)},\dots,f_{(m)}\}$ with $f_{(j)}$ the $j$-th largest among the $f(x_{i})$’s so that $f_{(j)}$ is the $(j/m)$ sample quantile of $Y$. Then, given a constant $\alpha\in[0,1]$, and denoted with $\lfloor j\rfloor$ the greatest integer less than or equal to $j$, we have that, in probability,
[…]
Basically, the HDR is derived based on the sample quantile of the density $f$. However, the density function itself is often unknown and we have to estimate it based on a set of available iid observations $s_{n}\doteq\{x_{1},\dots,x_{n}\}$. In this case, the $100(1-\alpha)\%$ HDR can be estimated as
$\hat{C}_{n}(\hat{f}_{\alpha})$ $\doteq C(x\!:f_{n}(x)>f_{\lfloor\alpha n\rfloor}),$ (1)
with $f_{n}$ being a possibly consistent estimator of $f$. Note that for small $n$, it may not be possible to get a reasonable density estimate. Also, with few observations and no prior knowledge on the underlying density function, there seems to be little point in attempting to summarize the density.
2.2 One-class neighbor machines (OCNM) approach
An alternative approach to HDR estimation, inspired by the theory of support vector machines (Schölkopf et al.,, 2001) and the notion of neighborhood measures, has been introduced in the machine learning literature by Munoz and Moguerza, (2006). The procedure is outlined in Proposition 2, after introducing neighborhood measures in Definition 2.
Definition 2 (Neighbourhood Measure; Munoz and Moguerza, (2006))
Let $X$ be a random variable with density function $f$ defined on $\mathbb{R}^{d}$. Denoted by $S_{n}$ the set of random iid samples $s_{n}=\{x_{1},\dots,x_{n}\}$ of size $n$ (drawn from $f$), the real-valued function $g\!:\mathbb{R}^{d}\times S_{n}\to\mathbb{R}$ is a neighbourhood measure if one of the following holds:
[…]
The function $g$ is called either a (a) sparsity or a (b) concentration measure.
Proposition 2 (Munoz and Moguerza, (2006))
Consider an iid sample $s_{n}=\{x_{1},\dots,x_{n}\}$ and a sparsity measure $g$. Define $\rho^{*}=g(x_{(\nu n)},s_{n})$, with $x_{(\nu n)}$ being the $(\nu n)$-th sample in the order induced in $s_{n}$ by $g$, provided that $\nu n\in\mathbb{N}$; otherwise, the least integer greater than $\nu n$, denoted by $\lceil\nu n\rceil$, is taken. Then, the binary decision function $h(x)=\operatorname{sign}(\rho^{*}-g(x,s_{n}))$ is such that:
[…]
where $\mathbb{I}$ denotes the indicator function, and
[…]
where $C(f_{\nu})$ is the minimum-volume set such that $\mathbb{P}(C(f_{\nu}))\geq\nu$.
Notably, the OCNM algorithm relaxes the HDR estimation problem in the following sense: Instead of estimating and evaluating the density $f$, a more general and potentially simpler measure $g$ that asymptotically preserves the order induced by the density, can be considered. It is, however, important to remark that the quality of the estimation procedure heavily depends on using a neighborhood measure. If the measure used is neither a concentration nor a sparsity measure, there is no reason why the method should work.
2.3 Neighborhood-quantile approach
We now discuss a hybrid approach that can be viewed as: i) a generalization of the density quantile approach and; ii) a restatement of the OCNM method directly in terms of its solution $\rho^{*}$. Compared to the former, it ensures a wider applicability allowing for a more general set of functions, including the density $f$ as a special case; with respect to the latter, it offers a more direct, interpretable, and computationally efficient method.
Theorem 1
Let $X$ be a continuous random variable defined on $\mathbb{R}^{d}$ with density function $f$, and consider a set $s_{n}=\{x_{1},\dots,x_{n}\}\in S_{n}$ of size $n$ (drawn from $f$). Assume $g\!:S_{n}\times\mathbb{R}^{d}\to\mathbb{R}$ is a neighbourhood measure. Then an estimate of the $100(1-\alpha)\%$ HDR can be obtained as:
[…]
where $g(x_{\lfloor\alpha n\rfloor},s_{n})$ is the $\alpha$-quantile of the sample $\{g(x_{1},s_{n}),\dots,g(x_{n},s_{n})\}$, and $\lfloor j\rfloor$ denotes the greatest integer less than or equal to $j$.
The proof is straightforward when considering the relationship between the region $C^{\text{OCNM}}_{n}$ as defined in Proposition 2 and the region $C_{n}$ as defined in Theorem 1. In fact, without loss of generality, taken $g$ to be a sparsity measure, and noticing that $\nu$ plays the role of $1-\alpha$, we can observe that:
$C^{\text{OCNM}}_{n}$ =$=\{x\!:h(x)=\operatorname{sign}(\rho^{*}-g(x,s_{n}))\geq 0\}$
=$=\{x\!:\operatorname{sign}(g(x_{\lceil\nu n\rceil},s_{n})-g(x,s_{n}))\geq 0\}$
=$=\{x\!:g(x,s_{n})\leq g(x_{\lceil(1-\alpha)n\rceil},s_{n})\}$
$=C_{n}.$
Remark If $g$ is chosen to be a concentration measure, then to ensure a $100(1-\alpha)\%$ coverage (notice that in this case $\mathbb{P}(C(f_{\nu}))=1-\nu$), the decision value $\rho^{*}$ induced by the concentration measure is given by $\rho^{*}=g(x_{\lfloor\alpha n\rfloor},s_{n})$.
Noticing that the density represents a concentration measure, the resemblance with the density quantile approach in Proposition 1 should now be clear. In fact, from the perspective of the density quantile approach, we can view Theorem 1 as a generalization of Proposition 1, where $f$ is replaced by any function $g$ that satisfies the criteria of neighborhood measures. Intuitively, provided that the density function $f$ is replaced by a function that preserves the order induced in the sample by $f$, the estimated HDR is asymptotically valid. Neighborhood measures ensure this ranking. To see it, without loss of generality, consider a sparsity measure $g$ and a sample $s_{n}=\{x_{1},\dots,x_{(1-\alpha)n},\dots,x_{n}\}$ ordered so that
$f(x_{1})<\dots<f(x_{(1-\alpha)n})<\dots<f(x_{n}),$
where, for simplicity, we suppose $(1-\alpha)n\in\mathbb{N}$ and $f(x_{j})\neq f(x_{j})$ for all $i\neq j$. From Definition 2 (a), for each pair $(x_{i},x_{j})$, with $i<j$, it holds that $\mathbb{P}\left(g(x_{i},s_{n})>g(x_{j},s_{n})\right)\to 1$. Hence, given $\epsilon\in(0,1)$, there exists $n_{ij}\in\mathbb{N}$ such that $\mathbb{P}\left(g(x_{i},s_{n_{ij}})>g(x_{j},s_{n_{ij}})\right)>1-\epsilon$. Taking $n\geq\max\{n_{ij}\}$, it is guaranteed that $\mathbb{P}\left(g(x_{1},s_{n})>\dots>g(x_{(1-\alpha)n},s_{n})>\dots>g(x_{n},s_{n})\right)>1-\epsilon$. Therefore, as $n\to\infty$
$\mathbb{P}\left(g(x_{1},s_{n})>\dots>g(x_{(1-\alpha)n},s_{n})>\dots>g(x_{n},s_{n})\right)\to 1.$
Remark Proposition 1 represents a special case of Theorem 1 with $g(x,s_{n})$ being either: (a) a concentration measure $g(x,s_{n})\propto\hat{f}(x,s_{n})$ or; (b) a sparsity measure $g(x,s_{n})\propto\frac{1}{\hat{f}(x,s_{n})}$, where $\hat{f}$ can be any consistent density estimator. Among the plethora of density estimators, in this work we focus on KDE (Parzen,, 1962), and use it as a benchmark measure for the proposed comparators in Section 3. Specifically, given a set of iid observations $s_{n}\doteq\{x_{1},\dots,x_{n}\}$ drawn from an unknown target density $f$, the KDE measure $M_{0}(x,s_{n},h)$ at the location $x$ is defined as
$M_{0}(x,s_{n},h)=\frac{1}{n}\sum_{i=1}^{n}K_{h}\left(\lVert x-x_{i}\rVert\right)=\frac{1}{nh^{d}}\sum_{i=1}^{n}K\left(\frac{\lVert x-x_{i}\rVert}{h}\right),$ (2)
[…]
3 Alternative Measures for Building HDRs
We now discuss a number of neighborhood measures that could be used to estimate HDRs. We repurpose some measures that have been successfully employed in areas such as classification or clustering, and introduce some novel ideas, included a copula-based approach.
3.1 $\boldsymbol{k}$NN-Euclidean distance
This corresponds to the sum of all Euclidean distances of a given point $x$ from its $k$-nearest neighbors. The concept of closeness (“nearest”) is defined according to the Euclidean metric or $L^{2}$-norm denoted by $\lVert\cdot\rVert_{2}$.
Definition 3 (kNN-Euclidean distance)
Given a data point $x\in\mathbb{R}^{d}$ of a sample set $s_{n}$ of size $n$, and an integer $k\in[1,n]$, we define the k-nearest neighborhood Euclidean distance of point $x$ from sample $s_{n}$ as:
$M_{1}(x,s_{n},k)\doteq\sum_{i=1}^{k}\lVert x-x_{(i)}\rVert_{2},$
where $x_{(i)}$ denotes the $i$-th observation in the reordered data such that $\lVert x-x_{(1)}\rVert_{2}\leq\dots\leq\lVert x-x_{(i)}\rVert_{2}\leq\dots\lVert x-x_{(n)}\rVert_{2}$.
$M_{1}(x,s_{n},k)$ is entirely based on a metric distance, that is, the Euclidean metric, and represents a sparsity measure. The property follows from the convergence in probability of the $k$-nearest neighbor density estimator (Silverman,, 1996). Similar distances have been used in the literature for nonparametric classification and regression, within the so-called $k$-nearest neighbors ($k$NN) algorithm. Notably, the distance to the nearest neighbors can also be seen as a local density estimate (Loftsgaarden and Quesenberry,, 1965; Silverman,, 1996) and as a special case of a variable-bandwidth KDE with a uniform kernel (Terrell and Scott,, 1992). A large $k$NN distance indicates that the density is usually small and vice versa. Furthermore, by ranking each data point according to the distance from its $k$-nearest neighbors, this measure can also be used as an outlier score in anomaly detection (Ramaswamy et al.,, 2000).
Compared to other methods, the $k$NN approach has several advantages such as: (i) being purely nonparametric, hence able to flexibly adapt to any continuous distribution; (ii) having a reasonable time complexity; (iii) depending on a unique hyperparameter $k$, whose tuning is relatively simple. The choice of $k$ should be made based on the sample data. Generally, higher values of $k$ reduce the effect of noise; however, they underfit the model–making boundaries between classes less distinct–and are computationally more expensive, especially for large $d$. A general rule of thumb in classification is $k=\lfloor\sqrt{n}\rfloor$, with $n$ the number of samples in the dataset. Further investigation of the role of $k$ in different settings and how it relates to the missclassification error is given in the Supplementary Material B and in (Meeker et al.,, 2017).
3.2 $\boldsymbol{k}$NN-CDF distance
Although the Euclidean distance and other general notions of measures (including concepts like area or volume) can certainly be useful for capturing relevant insights on the topology of a given set, they may be equipped with a probability measure to better resemble the notion of density. Let $\mathbb{P}$ be a probability measure and denote by $F$ its associated CDF and by $F_{i}$ the CDF of the $i$-th marginal.
Given two data points $x_{1}=(x_{11},\dots,x_{1d})\in\mathbb{R}^{d}$ and $x_{2}=(x_{21},\dots,x_{2d})\in\mathbb{R}^{d}$ from the sample set $s_{n}$, a preliminary version of the CDF distance between the two data points, denoted by $d_{\mathbb{P}}(x_{1},x_{2})$, has been defined in Venturini, (2015) in the context of clustering problems as:
$d_{\mathbb{P}}(x_{1},x_{2})=\sqrt{\sum_{j=1}^{d}\left(F_{j}(x_{1j})-F_{j}(x_{2j})\right)^{2}}.$ (3)
In the typical case of unknown marginal CDFs, their empirical counterpart $F_{n,j}(t)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(x_{ij}\leq t)$, $j=1,\dots,d$, could be considered.
The distance $d_{\mathbb{P}}$ can be interpreted as the composition of the (nonlinear) transformation $F\!:\mathbb{R}^{d}\to[0,1]$ and the computation of the ordinary Euclidean distance. It fulfills all the properties of being a proper metric. Furthermore, it has the nice property that the distance between two points is proportional to the probability contained between the two points. However, as defined in Eq. (3), it presents some key limitations for inferring the underlying density and estimating HDRs. Consider for simplicity the univariate Gaussian mixture model illustrated in Figure 3 and the set of points $x_{1}=-2$, $x_{2}=-1.6$, $x_{3}=-0.72$, $x_{4}=0.5$. When evaluating the CDF distance between $x_{2}$ and all other points, we have that $d_{\mathbb{P}}(x_{2},x_{1})\approx d_{\mathbb{P}}(x_{2},x_{3})$, but there is little to say about the density around $x_{1}$ and $x_{3}$, apart from realizing that they are clearly different. We could gain more information if we were to compute the CDF distance from the global mode $x_{1}$, assuming we have this information, noticing that the higher $d_{\mathbb{P}}(x_{1},x_{i})$, for all $i$, the lower the density of $x_{i}$. However, this is no longer verified for multimodal densities. In fact, we have $d_{\mathbb{P}}(x_{1},x_{4})>d_{\mathbb{P}}(x_{1},x_{3})$, but $f(x_{4})>f(x_{3})$, contrasting with the definition of a neighborhood measure.
Figure 3: Illustration of the CDF-distances computed on the set of points $x_{1}=-2$, $x_{2}=-1.6$, $x_{3}=-0.72$, $x_{4}=0.5$ based on a Gaussian mixture model.
Motivated by these limitations and inspired by the idea of the $k$NN distance in Definition 3, we propose a new variant consisting of the sum of CDF-distances between a point $x$ and its $k$-nearest neighbors. Here, “nearest” should again be understood according to the Euclidean distance. The rationale is simple: the higher the CDF distances between a point and its neighbors, the higher we expect the density to be for that point. Since the $k$ neighbors of each data point will be at different (Euclidean) distances compared to the $k$ neighbors of the other data points, we want to properly scale or weight the CDF distances according to this information. The proposed measure is reported in Definition 4.
Definition 4 ($\boldsymbol{k}$NN-CDF distance)
Given $F_{i}$, the $i$-th marginal of a CDF $F$, for $i=1,\dots,d$, and an integer $k\in[1,n]$, we define the $k$-nearest neighborhood CDF distance as:
$M_{2}(x,s_{n},k)=\begin{cases}\sum\limits_{i=2}^{k}\frac{d_{\mathbb{P}}(x,x_{(i)})}{\lVert x-x_{(i)}\rVert_{2}}=\sum\limits_{i=2}^{k}\frac{\sqrt{\sum_{j=1}^{d}\left(F_{j}(x_{j})-F_{j}(x_{(i)j})\right)^{2}}}{\sqrt{\sum_{j=1}^{d}\left(x_{j}-x_{(i)j}\right)^{2}}}&\quad k>1\\ 0&\quad k=1,\end{cases}$
where $x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}$, and $x_{(i)}=(x_{(i)1},\dots,x_{(i)d})$ is the $i$-th observation in the sample $s_{n}$ such that $\lVert x-x_{(1)}\rVert_{2}\leq\dots\leq\lVert x-x_{(i)}\rVert_{2}\leq\dots\lVert x-x_{(n)}\rVert_{2}$.
Property 1 (Semimetric)
Consider the $M_{2}$ measure as defined in Definition 4. It can be easily verified that, for any $k\in[1,n]$, $M_{2}(x,x_{(k)},k)\!:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ is a semimetric. In fact, for any set $x,x_{(k)}\in\mathbb{R}^{d}$, we have that:
  • If $x=x_{(k)}$, $M_{2}(x,x_{(k)},1)=0$ by definition.
  • If $x\neq x_{(k)}$, $M_{2}(x,x_{(k)},k)>0$. In fact, $(x_{j}-x_{(k)j})^{2}>0$ and $(F_{j}(x_{j})-F_{j}(x_{(k)j}))^{2}>0$ for all $j$ and all $k$, when $x\neq x_{(k)}$ and $X$ is continuous.
  • $M_{2}(x,x_{(k)},k)=M_{2}(x_{(k)},x,k)$, as $(x_{j}-x_{(k)j})^{2}=(x_{(k)j}-x_{j})^{2}$ and $(F_{j}(x_{j})-F_{j}(x_{(k)j}))^{2}=(F_{j}(x_{(k)j})-F_{j}(x_{j}))^{2}$ for all $j$ and all $k$.
The triangular inequality, stating that $M_{2}(x,x_{(k)},1)\leq M_{2}(x,x_{(k^{\prime})},1)+M_{2}(x_{(k^{\prime})},x_{(k)},1)$, for all $x,x_{(k)},x_{(k^{\prime})}\in\mathbb{R}^{d}$, does not hold. It is enough to consider a simple counterexample such as: $d=1$, $F$ the CDF of a standard normal distribution, and three points $x=1<x_{(k)}=2<x_{(k^{\prime})}=3$.
3.3 $\boldsymbol{\epsilon}$-neighborhood multivariate CDF distance
The measure $M_{2}(x,s_{n},k)$ introduced in Section 3.2 has several advantages; for example: (i) it takes into account the probabilistic information in the data, (ii) it is based on marginal CDFs which may be easy to estimate, and (iii) it has convenient computational times. However, using only the marginal CDFs may compromise their validity whenever the marginal components of a multivariate random variable $X\in\mathbb{R}^{d}$ show a dependence structure. In that case, the use of a multivariate CDF may be more appropriate. If the multivariate CDF is unknown, we could use its empirical counterpart. Notice, however, that this estimation part may compromise computational efficiency, especially in high dimensions and for large values of $k$. We therefore focus on an $\epsilon$-neighborhood version, which reduces the number of operations from $k$ to $1$, once $\epsilon$ is determined.
Definition 5 ($\epsilon$-neighborhood multivariate CDF distance)
Given the CDF $F$ of a multivariate variable $X\in\mathbb{R}^{d}$, we define the $\epsilon$-neighborhood multivariate CDF distance of a point $x=(x_{1},\dots,x_{d})$ to be the probability of that point belonging to a hyperrectangle of dimension $d$ defined on vertexes $[x_{1}-\epsilon_{1},x_{1}+\epsilon_{1}]\times\dots\times[x_{d}-\epsilon_{d},x_{d}+\epsilon_{d}]$ scaled by hyperrectangle’s $d$-volume:
$M_{3}(x,\epsilon)$ $=\frac{\mathbb{P}(X\in[x-\epsilon,x+\epsilon])}{\prod_{j=1}^{d}2\epsilon_{j}}=\frac{\sum_{\nu\in\mathcal{V}}(-1)^{n(\nu)}F(\nu)}{\prod_{j=1}^{d}2\epsilon_{j}}\propto\sum_{\nu\in\mathcal{V}}(-1)^{n(\nu)}F(\nu),$ (4)
where $\nu=(\nu_{1},\dots,\nu_{d})$, with $\nu_{j}\in\{x_{j}-\epsilon_{j},x_{j}+\epsilon_{j}\}$, for $j\in 1,\dots,d$, and $n(\nu)=\sum_{j=1}^{d}\mathbb{I}(\nu_{j}=x_{j}-\epsilon_{j})$. The sum is computed over the $2^{d}$ vectors of the set $\mathcal{V}$.
As in the case of the density $f$ itself, when the CDF $F$ is unknown, a consistent estimator, say $F_{n}$, may be used leading to:
$M_{3}(x,s_{n},\epsilon)$ $=\frac{\sum_{\nu\in\mathcal{V}}(-1)^{n(\nu)}F_{n}(\nu)}{\prod_{j=1}^{d}2\epsilon_{j}}\propto\sum_{\nu\in\mathcal{V}}(-1)^{n(\nu)}F_{n}(\nu).$
Relationship in Eq. (4) can be derived by recursion starting from small $d$ values. For $d=2$, for example, it is easy to verify that
$M_{3}(x,\epsilon)$ $\propto\mathbb{P}(X\in[x-\epsilon,x+\epsilon])$
$=\mathbb{P}(X_{1}\in[x_{1}-\epsilon_{1},x_{1}+\epsilon_{1}],X_{2}\in[x_{2}-\epsilon_{2},x_{2}+\epsilon_{2}])$
$=\mathbb{P}(X_{1}\in[x_{1}-\epsilon_{1},x_{1}+\epsilon_{1}],X_{2}\leq x_{2}+\epsilon_{2})-\mathbb{P}(X_{1}\in[x_{1}-\epsilon_{1},x_{1}+\epsilon_{1}],X_{2}<x_{2}-\epsilon_{2})$
$=\mathbb{P}(X_{1}\leq x_{1}+\epsilon_{1},X_{2}\leq x_{2}+\epsilon_{2})-\mathbb{P}(X_{1}<x_{1}-\epsilon_{1},X_{2}\leq x_{2}+\epsilon_{2})$
$\quad-\mathbb{P}(X_{1}\leq x_{1}+\epsilon_{1},X_{2}\leq x_{2}-\epsilon_{2})-\mathbb{P}(X_{1}<x_{1}-\epsilon_{1},X_{2}<x_{2}-\epsilon_{2})$
$=F(x_{1}+\epsilon_{1},x_{2}+\epsilon_{2})-F(x_{1}-\epsilon_{1},x_{2}+\epsilon_{2})-F(x_{1}+\epsilon_{1},x_{2}-\epsilon_{2})+F(x_{1}-\epsilon_{1},x_{2}-\epsilon_{2})$
$=\sum_{\nu\in\mathcal{V}}(-1)^{n(\nu)}F(\nu),$
with $\mathcal{V}=\{(x_{1}+\epsilon_{1},x_{2}+\epsilon_{2}),(x_{1}-\epsilon_{1},x_{2}+\epsilon_{2}),(x_{1}+\epsilon_{1},x_{2}-\epsilon_{2}),(x_{1}-\epsilon_{1},x_{2}-\epsilon_{2})\}$.
We now state the following result for this measure.
Property 2 (Density equivalence)
Consider the $M_{3}$ measure as defined in Definition 5 and assume that $F$ is differentiable at any point $x$ of its support. Then, as $\epsilon\to 0$, $M_{3}(x,\epsilon)\to f(x)$, for any $x\in\mathbb{R}^{d}$.
The proof follows from basic probability and calculus theory.
3.4 Copula-based measures
A $d$-dimensional copula $C\!:[0,1]^{d}\to[0,1]$ is a CDF with uniform marginal distribution functions (Nelsen,, 2006). If we consider a random vector $X=(X_{1},\dots,X_{d})$ with joint CDF $F$ and marginals $F_{1},\dots,F_{d}$, then the copula of $X$ is represented by the joint distribution of $F_{1}({X_{1}}),\dots,F_{d}({X_{d}})$ and is derived as:
$C(u_{1},\dots,u_{d})$ $=\mathbb{P}(F_{1}({X_{1}})\leq u_{1},\dots,F_{d}({X_{d}})\leq u_{d})$
$=\mathbb{P}(X_{1}\leq F_{1}^{-1}(u_{1}),\dots,X_{d}\leq F_{d}^{-1}(u_{d}))$
$=F(F_{1}^{-1}(u_{1}),\dots,F_{d}^{-1}(u_{d})).$
Letting $u_{j}\doteq F_{j}(x_{j})$, this yields the following well-known result due to Sklar, (1959):
$F(x_{1},\dots,x_{d})=C(F_{1}(x_{1}),\dots,F_{d}(x_{d})),\quad\text{for all}\ x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}.$ (5)
Furthermore, when the random vector $X$ is continuous with density $f$, we also have that
$f(x_{1},\dots,x_{d})=c\Big{(}F_{1}(x_{1}),\dots,F_{d}(x_{d})\Big{)}\times f_{1}(x_{1})\times\cdots\times f_{d}(x_{d}),$ (6)
where $c$ is the density of the random vector $\Big{(}F_{1}(x_{1}),\dots,F_{d}(x_{d})\Big{)}\in[0,1]^{d}$.
In summary, we can decompose every $d$-dimensional CDF $F$ or PDF $f$ into a composition of their marginal distribution functions and a $d$-copula. This allows us to redefine both $M_{0}$ in Eq. (2) and $M_{3}$ in Definition 5 in terms of their copula representation. For example, in the case of the $M_{3}$ measure, we may construct its copula-based alternative given in Definition 6.
Definition 6 ($\epsilon$-neighborhood copula-based CDF distance)
Given the CDF $F$ of a random vector $X=(X_{1},\dots,X_{d})$, we define the $\epsilon$-neighborhood copula-based CDF distance of point $x=(x_{1},\dots,x_{d})$ to be
[…]
where $\nu=(\nu_{1},\dots,\nu_{d})$, with $\nu_{j}\in\{x_{j}-\epsilon_{j},x_{j}+\epsilon_{j}\}$, for $j\in 1,\dots,d$, and $n(\nu)=\sum_{j=1}^{d}\mathbb{I}(\nu_{j}=x_{j}-\epsilon_{j})$.
For $d=2$, it is straightforward to see that
$M^{\text{Cop}}_{3}(x,\epsilon)$ $\propto C\left(F_{1}(x_{1}+\epsilon_{1}),F_{2}(x_{2}+\epsilon_{2})\right)-C(F_{1}(x_{1}-\epsilon_{1}),F_{2}(x_{2}+\epsilon_{2}))$
$\quad-C(F_{1}(x_{1}+\epsilon_{1}),F_{2}(x_{2}-\epsilon_{2}))+C(F_{1}(x_{1}-\epsilon_{1}),F_{2}(x_{2}-\epsilon_{2})).$
The main advantage of this representation over the one involving the joint CDF is that the estimation of the multivariate distribution–when unknown–is performed through the estimation of the (univariate) marginals, evading thus the curse of dimensionality (see e.g., Nagler and Czado,, 2016). Furthermore, copulae offer a flexible framework that captures complex dependency structures while offering direct control over the marginals.
4 Empirical Evaluation
4.1 Simulation setup
We now evaluate and compare the proposed measures in an extensive number of simulated bivariate settings that vary according to the complexity of the data, as defined below.
  1. (i)
    We consider different copula models for the dependence structure among marginals, all sharing the same degree of dependence, expressed via Kendall’s $\tau$, and set equal to $0.5$. The list of different copulae follows.
    $C^{\text{Gauss}}_{\theta=0.7}$
    Gaussian family (Elliptical class) with zero tail dependence and radial symmetry.
    $C^{\text{t}}_{\theta=0.7,\nu=6}$
    Student-t family (Elliptical class) with radial symmetry.
    $C^{\text{Frank}}_{\theta=5.75}$
    Frank family (Archimedian class) with radial symmetry.
    $C^{\text{Clay}}_{\theta=2}$
    Clayton family (Archimedian class) not restricted to radial symmetry.
  2. (ii)
    Each of the above copulae is then combined with different simulation schemes for the marginals, according to the following details. Fixing $\sigma^{2}=2$ and $w_{1}=1-w_{2}=0.5$, and taking $\mu_{11}=0,\mu_{12}=9,\mu_{21}=1,\mu_{22}=8$, we consider the following.
    Unimodal–heavy tails
    Student-t model $X_{i}\sim t_{\nu=2},\quad i=1,2$.
    Unimodal
    Gaussian model $X_{i}\sim\mathcal{N}(\mu_{i1},\sigma^{2}),\quad i=1,2$.
    Bimodal
    Gaussian $X_{1}\sim\mathcal{N}(\mu_{11},\sigma_{1}^{2})$ Gaussian mixture $X_{2}\sim\sum_{k=1}^{2}w_{k}\mathcal{N}(\mu_{2k}+5\mathbb{I}(k=2),\sigma^{2})$.
    Quadrimodal
    Gaussian mixture $X_{i}\sim\sum_{k=1}^{2}w_{k}\mathcal{N}(\mu_{ik},\sigma^{2}),\quad i=1,2$.
  3. (iii)
    We also consider a distribution on a compact supports, that is, the Dirichlet distribution defined on the $(K-1)$-simplex, with $K=3$ and with parameters $\mathbf{\alpha}=(1,1,2)$. In this case, the dependence structure is uniquely determined and its closed-form copula expression is given in Section 4.4.1.
  4. (iv)
    We also vary the sample size, with $n\in\{50,100,500,1000\}$.
In total, 17 different scenarios are considered, named S1 to S17. Their graphical representation, in the form of contour plots, is provided in Figure 4 (S1-S16) and Figure 5 (S17).
Figure 4: Evaluated scenarios with their contour plots at different levels, included the one delimiting the $95\%$ HDR (blue color).
4.2 Evaluated measures
In all settings, we employ the proposed neighborhood-quantile method (see Section 2.3) on data samples of different sizes generated according to the aforementioned scenarios. For each scenario, we evaluate the following eight methods, with full details on their hyperparameter tuning reported in the Supplementary Material B.
$M_{0}$:KDE
Direct estimation of the bivariate density. We use KDE, with Gaussian kernel and bandwidth selection based on the asympotically optimal solution proposed in Chacón et al., (2011), where its adequacy is shown in general settings, including Gaussian mixture models.
$M^{\text{NPCop}}_{0}$:DE
Nonparametric indirect density estimation with copula. We use standard KDE with the same optimal bandwidth of Chacón et al., (2011) for the univariate marginals, and KDE with the transformation local likelihood estimator and nearest-neighbor bandwidth for the copula density. We refer to Nagler and Czado, (2016) for details.
$M^{\text{PCop}}_{0}$:DE
Parametric indirect density estimation with copula. We adopt a fully parametric approach (with maximum likelihood fitting) to estimate both marginals and the copula. For the copula model, we select the best model using the AIC criterion; no misspecification is introduced for the marginals.
$M_{1}$:$k$NN-Eucl
Cumulative Euclidean distances from the $k$NNs. We sum the Euclidean distances between each point and its $k$ neighbors defined according to the Euclidean metric. The choice of the hyperparameter $k$ is based on an extensive cross-validation procedure, leading to a general rule of thumb aligned with the existing literature: $k=\left[\sqrt{n/2}\right]$, with $\left[x\right]$ the integer closest to $x$.
$M_{2}$:$k$NN-CDF
Cumulative CDF distances from the $k$NNs. We sum the CDF distances between each point and its $k$ neighbors, which are defined according to the Euclidean metric. The measure follows a univariate approach as detailed in Section 3.2. The choice of the hyperparameter $k$ is based on an extensive cross-validation procedure, suggesting a uniform choice across different scenarios and sample sizes: $k=30$.
$M_{3}$:$\epsilon$-CDF
This represents the $\epsilon$-neighborhood multivariate CDF distance introduced in Section 3.3. The empirical CDF is used as the CDF estimator. The optimal choice of the hyperparameter $\epsilon$ is based on an extensive cross-validation procedure, leading to the following heuristic for S1-S16 as a result of an exponential decay model fit with respect to the sample size: $\epsilon=\exp\left(2.13-0.3\log(n)\right)$.
$M^{\text{NPCop}}_{3}$:$\epsilon$-CDF
Fully nonparametric indirect estimation of the $\epsilon$-CDF measure with copula. We use the empirical CDF for estimating the univariate marginals, and KDE with the transformation local likelihood estimator and nearest-neighbor bandwidth for the copula density. The optimal choice of the hyperparameter $\epsilon$ follows the same strategy as the previous measure, leading to the following heuristic for S1-S16: $\epsilon=\exp\left(1.74-0.26\log(n)\right)$.
$M^{\text{PCop}}_{3}$:$\epsilon$-CDF
Parametric indirect CDF estimation with copula. We adopt a fully parametric approach (with maximum likelihood fitting) to estimate both marginals and the copula. For the copula model, we select the best model using the AIC criterion; no misspecification is introduced for the marginals. The optimal choice of the hyperparameter $\epsilon$ follows the heuristic $\epsilon=\exp\left(1.60-0.41\log(n)\right)$ for S1-S16.
Exception made for the measures depending on $\epsilon$ (with considerations deferred to Section 4.4.1), the same hyperparameter choices are adopted in S17.
4.3 Performance metrics
The measurement of the performance of an HDR estimator can be closely related to the specific problem of interest. In this work, for example, our main interest is in detecting abnormal values that, under an HDR framework, would correspond to those points that fall outside the region. This motivates certain metrics of common use in one-class classification problems, which we also employ in this work. However, we emphasize that the derivation of an HDR can in principle have a wider scope compared to a classification goal. Indeed, it would define the region of highest-density points (e.g., normal values), regardless of whether these points have been observed or not. Thus, it would not only allow us to classify an observed point as normal or abnormal, but it would also provide the entire region of normal values, useful, for example, in a prediction setting prior to observing a point.
With that being said, let FP, TP, FN, and TN be the number of false positive, true positive, false negative, and true negative points, respectively, where positive refer to those points that should be outside the true $(1-\alpha)\%$ HDR and negative the others:
$TN=\sum_{i\in s_{n}}\mathbb{I}(x_{i}\in C(f_{\alpha})),\quad\quad FN=\sum_{i\in s_{n}}\mathbb{I}(x_{i}\in\hat{C}_{n}(\hat{f}_{\alpha})\mid x_{i}\notin C(f_{\alpha})),$
$TP=\sum_{i\in s_{n}}\mathbb{I}(x_{i}\notin C(f_{\alpha})),\quad\quad FP=\sum_{i\in s_{n}}\mathbb{I}(x_{i}\notin\hat{C}_{n}(\hat{f}_{\alpha})\mid x_{i}\in C(f_{\alpha})).$
Well-established measures of inefficiency are false negative/positive rates (FNR/FPR), and the total error rate (ERR), that is, the one-complement of accuracy:
$\text{FNR}=\frac{FN}{FN+TP},\quad\text{FPR}=\frac{FP}{FP+TN},\quad\text{ERR}=\frac{FN+FP}{FN+FP+TN+TP}=1-\text{Accuracy}.$
To account for the potentially high imbalance between positives and negatives, we also evaluate the two-sided F1 score, and the Matthews correlation coefficient (MCC; Matthews,, 1975) alternatively known in statistics as the $\phi$-coefficient (see pg. 282 in Cramér,, 1946):
F1 $=\frac{2TP}{2TP+FP+FN}+\frac{2TN}{2TN+FP+FN},$
MCC $=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)\times(TP+FN)\times(TN+FP)\times(TN+FN)}}.$
In particular, MCC has been shown to produce good scores only if the classification is adequate in all four elements of interest (true positives, false negatives, true negatives, and false positives), overcoming the overoptimistic inflated results, especially on imbalanced datasets, of other popular classification measures (Chicco and Jurman,, 2020). All evaluations are based on $\alpha=0.05$, that is, a coverage probability of 95%, or, alternatively stated, $5\%$ and $95\%$ of positives and negatives, respectively.
4.4 Simulation results
Comparative performance results are reported in Table 1. Due to space limitation, we focus on four scenarios that are representative of the different copula-induced dependencies and multimodalities, specifically those on the diagonal of Figure 4 (S1, S6, S11, S16), deferring to the Supplementary Material C all the scenarios and sample sizes considered. The Dirichlet case is covered in Section 4.4.1.
In general, copula-based approaches result in the most performing measures, with classification errors (ERR, FPR, FNR) uniformly smaller–or at most equal–than those of the other measures and at no cost in terms of variability. Their advantage is particularly interesting in more complex scenarios, such as the quadrimodal case, where the parametric copula approach (both $M^{\text{PCop}}_{0}$:DE and $M^{\text{PCop}}_{3}$:$\epsilon$-CDF) practically halves the total error rate (ERR) of a standard $M_{0}$:KDE. Although this distance may be considered relatively negligible when looking at the ERR (maximum distance of $0.015$ in S16) and the FPR (maximum distance of $0.008$ in S16), it plays a significant role for the FNR, with an error difference of $0.157$. In practice, the probability of a true positive being missed by the “test” or measure decreases from around $28\%$ ($M_{0}$:KDE; S16) to around $13\%$ ($M^{\text{PCop}}_{0}$:DE, $M^{\text{PCop}}_{3}$:$\epsilon$-CDF; S16). This aspect is well-captured by the alternative performance metrics of F1 and MCC, which offer a more reliable global measure of efficiency compared to the ERR or the Accuracy. As depicted in Table 1–S16, when comparing $M_{0}$:KDE to $M^{\text{PCop}}_{0}$:DE or $M^{\text{PCop}}_{3}$:$\epsilon$-CDF, the MCC, for example, increases from $0.693$ to more than $0.85$. Note that MCC varies from $-1$ (worst value) to $1$ (best value).
When comparing the measures without the copula, no substantial differences can be noticed between $M_{0}$:KDE, $M_{1}$:$k$NN-Eucl and $M_{3}$:$\epsilon$-CDF, with a slightly improved performance of the second. The worst behavior is shown by $M_{2}$:$k$NN-CDF (with FNR being as high as $72\%$ in S16). Note that this was expected as $M_{2}$:$k$NN-CDF follows a univariate rationale, motivating therefore the multivariate $M_{3}$:$\epsilon$-CDF proposal (see Section 3).
Table 1: Performance results in terms of mean (standard deviation) of the compared methods in selected scenarios for sample size $n=500$.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
Scenario S1 (Unimodal): Gaussian copula – Student-t marginals
ERR 0.015 (0.006) 0.013 (0.006) 0.018 (0.007) 0.017 (0.006) 0.015 (0.007) 0.009 (0.006) 0.014 (0.006) 0.009 (0.006)
FPR 0.008 (0.005) 0.007 (0.006) 0.009 (0.007) 0.010 (0.006) 0.008 (0.006) 0.004 (0.005) 0.007 (0.006) 0.005 (0.006)
FNR 0.135 (0.092) 0.117 (0.079) 0.173 (0.075) 0.132 (0.089) 0.139 (0.088) 0.073 (0.087) 0.125 (0.086) 0.075 (0.086)
Accuracy 0.985 (0.006) 0.987 (0.006) 0.982 (0.007) 0.983 (0.006) 0.985 (0.007) 0.991 (0.006) 0.986 (0.006) 0.991 (0.006)
F1 1.842 (0.061) 1.863 (0.063) 1.807 (0.080) 1.826 (0.066) 1.840 (0.075) 1.908 (0.061) 1.854 (0.062) 1.906 (0.061)
MCC 0.846 (0.058) 0.867 (0.059) 0.811 (0.077) 0.830 (0.063) 0.844 (0.072) 0.912 (0.055) 0.858 (0.058) 0.910 (0.055)
Scenario S6 (Unimodal): Student-t copula – Gaussian marginals
ERR 0.015 (0.006) 0.017 (0.006) 0.060 (0.009) 0.019 (0.006) 0.014 (0.006) 0.011 (0.005) 0.017 (0.006) 0.011 (0.005)
FPR 0.008 (0.005) 0.009 (0.005) 0.032 (0.006) 0.010 (0.006) 0.007 (0.005) 0.006 (0.005) 0.009 (0.005) 0.006 (0.005)
FNR 0.140 (0.089) 0.161 (0.094) 0.601 (0.090) 0.165 (0.092) 0.128 (0.090) 0.096 (0.085) 0.157 (0.093) 0.095 (0.085)
Accuracy 0.985 (0.006) 0.983 (0.006) 0.940 (0.009) 0.981 (0.006) 0.986 (0.006) 0.989 (0.005) 0.983 (0.006) 0.989 (0.005)
F1 1.838 (0.062) 1.816 (0.065) 1.363 (0.096) 1.804 (0.068) 1.851 (0.059) 1.884 (0.058) 1.820 (0.061) 1.886 (0.058)
MCC 0.843 (0.059) 0.820 (0.062) 0.365 (0.096) 0.808 (0.066) 0.855 (0.055) 0.888 (0.053) 0.825 (0.059) 0.890 (0.053)
Scenario S11 (Bimodal): Frank copula – Gaussian & Gaussian mixture marginals
ERR 0.019 (0.007) 0.016 (0.006) 0.061 (0.009) 0.018 (0.006) 0.015 (0.006) 0.011 (0.006) 0.015 (0.006) 0.011 (0.006)
FPR 0.010 (0.005) 0.008 (0.005) 0.032 (0.006) 0.011 (0.006) 0.008 (0.005) 0.006 (0.005) 0.008 (0.005) 0.006 (0.005)
FNR 0.175 (0.095) 0.144 (0.093) 0.608 (0.092) 0.150 (0.093) 0.139 (0.093) 0.096 (0.089) 0.136 (0.093) 0.096 (0.089)
Accuracy 0.981 (0.007) 0.984 (0.006) 0.939 (0.009) 0.982 (0.006) 0.985 (0.006) 0.989 (0.006) 0.985 (0.006) 0.989 (0.006)
F1 1.800 (0.068) 1.832 (0.064) 1.355 (0.097) 1.810 (0.065) 1.837 (0.061) 1.882 (0.059) 1.841 (0.061) 1.883 (0.059)
MCC 0.804 (0.066) 0.836 (0.061) 0.357 (0.097) 0.815 (0.063) 0.842 (0.058) 0.887 (0.055) 0.845 (0.058) 0.887 (0.054)
Scenario S16 (Quadrimodal): Clayton copula – Gaussian mixture marginals
ERR 0.029 (0.007) 0.027 (0.008) 0.072 (0.010) 0.032 (0.008) 0.021 (0.007) 0.014 (0.006) 0.024 (0.007) 0.014 (0.006)
FPR 0.015 (0.006) 0.014 (0.006) 0.038 (0.005) 0.019 (0.007) 0.011 (0.006) 0.007 (0.006) 0.013 (0.006) 0.008 (0.006)
FNR 0.283 (0.092) 0.256 (0.093) 0.720 (0.088) 0.276 (0.095) 0.197 (0.089) 0.126 (0.089) 0.226 (0.088) 0.129 (0.089)
Accuracy 0.971 (0.007) 0.973 (0.008) 0.928 (0.010) 0.968 (0.008) 0.979 (0.007) 0.986 (0.006) 0.976 (0.007) 0.986 (0.006)
F1 1.689 (0.080) 1.716 (0.081) 1.238 (0.089) 1.670 (0.083) 1.777 (0.074) 1.850 (0.067) 1.748 (0.077) 1.848 (0.069)
MCC 0.693 (0.079) 0.720 (0.080) 0.239 (0.090) 0.675 (0.082) 0.782 (0.072) 0.855 (0.064) 0.752 (0.075) 0.852 (0.066)
4.4.1 Simplex scenario: Dirichlet
A particular interest is devoted to the simplex scenario. We consider a Dirichlet random vector $X\sim\mathcal{D}(\alpha_{1},\alpha_{2},\alpha_{3})$ defined on the $2$-dimensional simplex, where $\alpha_{j}>0,\ j=1,2,3$; we refer to Chapter XI in Devroye, (1986) for a detailed exposition. Up to a normalizing constant, the density of $X$ is given by
$f(x_{1},x_{2};\alpha_{1},\alpha_{2},\alpha_{3})\propto x_{1}^{\alpha_{1}-1}x_{2}^{\alpha_{2}-1}(1-x_{1}-x_{2})^{\alpha_{3}-1},\quad x_{1},x_{2}\in[0,1];x_{1}+x_{2}\leq 1.$
In the $2$-dimensional simplex, when $\alpha_{1}=1,\alpha_{2}=1,\alpha_{3}=a$, the associated copula density has the following expression up to a normalizing constant:
$c\left(u_{1},u_{2};\alpha_{1}=1,\alpha_{2}=1,\alpha_{3}=a\right)\propto\frac{\left[(1-u_{1})^{\frac{1}{a+1}}+(1-u_{2})^{\frac{1}{a+1}}-1\right]^{a-1}}{\left[(1-u_{1})(1-u_{2})\right]^{\frac{a}{a+1}}},$
with $u_{1},u_{2}$ such that $u_{1},u_{2}\in[0,1]$ and $(1-u_{1})^{\frac{1}{a+1}}+(1-u_{2})^{\frac{1}{a+1}}\geq 1$. The analytical derivation is given in Supplementary Material A, while its graphical representation–in terms of a set of random draws and different level sets–can be depicted in Figure 5.
Figure 5: Contours of a Dirichlet(1,1,2) density (left) and of its copula density (right). The $95\%$ HDR is depicted in blue color.
To evaluate the proposed measures, we first need to check or perform an accurate tuning of their hyperparameters. As shown in the Supplementary Material B, the optimal choice of $k$–which represents the optimal number of neighbors to take into account for computing both $M_{1}$:$k$NN-Eucl and $M_{2}$:$k$NN-CDF–remains the same as for the other scenarios: $k=\left[\sqrt{n/2}\right]$. However, for measures depending on $\epsilon$–which defines the length, area, or volume of the neighborhood–the optimal choice depends on the support of the underlying variable. Being defined on a simplex, in a Dirichlet scenario, the optimal values of $\epsilon$ have a smaller magnitude compared to noncompact or less restrictive scenarios such as S1-S16. In this case, the following choices are made for $\epsilon$, guided by simulation studies reported in the Supplementary Material B:
$M_{3}$:$\epsilon$-CDF
Performances are robust to sample size, with optimal $\epsilon=0.10$.
$M^{\text{NPCop}}_{3}$:$\epsilon$-CDF
Performances depend on the sample size, according to a heuristic given by: $\epsilon=\exp\left(-1.22-0.23\log(n)\right)$. This relationship is obtained by fitting a nonlinear regression (exponential decay model), with the empirical optimal $\epsilon$ and the sample size $n$ as dependent and independent variables, respectively.
$M^{\text{PCop}}_{3}$:$\epsilon$-CDF
Performances are robust to sample size, with optimal $\epsilon=0.02$.
In terms of results, as suggested in Table 1, the more complex the scenario, the more difficult we expect it to be for identifying the highest-density points (or true negatives) vs. the true positives. In this particular case, as shown in Figure 6, the discrepancy between the different measures is even more evident with high and persistently high FNR values, even with increasing sample sizes, for $M_{0}$:KDE and $M_{2}$:$k$NN-CDF. All other measures improve with the sample size, achieving results comparable to the other scenarios.
Figure 6: Performance results in terms of mean and error bounds of the compared methods in the Dirichlet scenario for varying sample sizes.
4.5 MAGIC data
We now apply the proposed measures to construct an HDR for the joint distribution of two selected variables from the MAGIC dataset (Bock et al.,, 2004). These data simulate the registration of high-energy gamma particles in a ground-based atmospheric Cherenkov gamma telescope, and have been studied in classification problems (Dvořák and Savický,, 2007), as well as to analyze the dependency structure of some of the variables (see e.g., Nagler and Czado,, 2016; Grazian et al.,, 2022).
In this evaluation, we focus on gamma-ray observations (overall $n=12,332$) and consider the two variables “fConc1” and “fM3Long”, after scaling them. We refer to Bock et al., (2004) for a full description of the dataset. In this case (as deduced from the complex structure of the data; see Figure 7), the parametric approach is inappropriate for both the estimation of the marginal distribution and, more importantly, the copula model. Thus, in Figure 7, we illustrate the derived 95% HDR using the nonparametric measures only. Although in the absence of the underlying truth it is not possible to perform a reliable evaluation, it seems that the two nonparametric copula-based approaches ($M^{\text{NPCop}}_{0}$:DE, $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF), as well as the distance-based $M_{1}$:$k$NN-Eucl and $M_{3}$:$\epsilon$-CDF, more sensibly exclude tail data points (which may be expected to have a lower density) from the HDR.
Figure 7: Estimated 95% HDR of two scaled variables (“fConc1” on x-axis and “fM3Long” on y-axis) of the MAGIC dataset. Only nonparametric measures are evaluated. Cadet-blue points define the estimated 95% HDR; in contrast, purple points are those lying outside the HDR.
To deal with estimation and data uncertainty, we also explore the potential of a “measure averaging” in a similar, but simplified, fashion to model averaging (see e.g., Hoeting et al.,, 1999; Hjort and Claeskens,, 2003). By forming a consensus between the different available measures, we expect the resulting HDR to be more robust than the individual estimates. Specifically, in Figure 8, we illustrate a 95% HDR formed by averaging across the different nonparametric measures: the estimated HDR is defined by the set of points identified as highest-density points by more than half of the evaluated metrics. Notably, when comparing the resulted HDR in Figure 8 and those obtained by the individual measures (Figure 7), we identify $M_{3}$:$\epsilon$-CDF and $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF as the closest regions (both with < 1% classification difference). Therefore, these may result in a possible superior ability to detect anomalous values.
Figure 8: Estimated 95% HDR based on a “measure averaging” approach of two scaled variables (“fConc1” on x-axis and “fM3Long” on y-axis) of the MAGIC dataset. Cadet-blue points define the estimated 95% HDR; in contrast, purple points are those lying outside the HDR.
5 Discussion and Conclusion
In this work, we have discussed some generalizations of the standard density-based approach to compute HDRs by using neighborhood measures. Various measures, with distinct properties, have been introduced and compared to the classical kernel density estimator, across different scenarios and sample sizes. Furthermore, the use of copula models for representing a multivariate distribution through a combination of its marginals and their dependence structure has been investigated. Our results suggest that such a generalized approach may provide great advantages to HDR estimation when considering several alternative measures to KDE, in particular those based on copulae (both under a parametric and nonparametric approach). In fact, compared to traditional KDE, the copula-based HDRs resulted in greater accuracy and lower FPR and FNR in a number of simulation scenarios and possibly in real data (as suggested by the MAGIC data application). Such performances are particularly important when the interest is in balancing different types of errors, minimizing both false positives and false negatives, and maximizing therefore the probability of detecting outliers or anomalous values. Among the different evaluated measures, those based on the $\epsilon$-neighborhood multivariate CDF distance ($M_{3}$:$\epsilon$-CDF, $M^{\text{PCop}}_{3}$:$\epsilon$-CDF, $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF) show the most promising results, especially in the nonparametric setting and when the copula is involved. Since the main characteristic of these measures is that they are based on the CDF rather than the PDF, it may suggest that focusing on estimating the former may be advantageous. Similar findings are reported by Magdon-Ismail and Atiya, (2002), who acknowledge that, compared to directly estimating the PDF with e.g., KDE, approximating the CDF is less sensitive to statistical fluctuations and its convergence rate is faster than the convergence rate of KDE methods. Although heuristic considerations for each specific measure are made in terms of their hyperparameters, e.g., the choice of $k$ in the kNN-based approaches, further work may investigate the existence of optimal theoretical values in a similar fashion to the plethora of asymptotic works on the optimal bandwidth choice in KDE (see e.g., Wand and Jones, 1994b, ; Chacón et al.,, 2011).
In this work, we focused on estimating HDR for continuous distributions that are dominated by the Lebesgue measure. Furthermore, despite providing a general multidimensional framework for building HDRs, we evaluated the proposed approach in a bivariate context. Future lines of research may examine the problem in the underappreciated setting of discrete probability distributions and in higher dimensions. In the first case, possible connections could be made with the work of O’Neill, (2022), which established some theory and an algorithm for HDRs in the space of discrete distributions. In the second case, and more specifically with reference to copula-based measures, our aim is to explore the use of vine copulae (Nagler and Czado,, 2016) to construct flexible dependence models for an arbitrary number of variables using only bivariate building blocks. We expect, in fact, to see remarkable advantages in using copulae over an increased number of variables, as the extension of the common KDE to high dimensions has proven challenging in terms of both computational efficiency and statistical inference.
Supporting information
Additional supporting information is provided in the Supplementary material, made available at the end of this document. The implementation of the algorithms and R codes to reproduce all the results and Figures/Tables of the main manuscript are provided at https://github.com/nina-DL/HDR_neigh_measures.
Supplementary Material
Appendix A Derivation of the Dirichlet copula
In this section, we provide the analytical derivation of the Dirichlet copula in the specific case of a $K-1$-dimensional simplex distribution, with $K=3$. Let $X$ be a Dirichlet random vector, denoted by $X\sim\mathcal{D}(\boldsymbol{\alpha})$, where $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3})$ and $\alpha_{j}>0,\ j=1,2,3$. Its joint density is given by
$f(x_{1},x_{2};\boldsymbol{\alpha})=\frac{1}{B(\boldsymbol{\alpha})}x_{1}^{\alpha_{1}-1}x_{2}^{\alpha_{2}-1}(1-x_{1}-x_{2})^{\alpha_{3}-1},\quad x_{1},x_{2}\in[0,1];x_{1}+x_{2}\leq 1,$
where $B(\cdot)$ is the multivariate Beta function defined in terms of the Gamma function $\Gamma(\cdot)$ as:
$B(\boldsymbol{\alpha})=\frac{\prod_{j=1}^{K}\Gamma(\alpha_{j})}{\Gamma\left(\sum_{j=1}^{K}\alpha_{j}\right)}=\frac{\Gamma(\alpha_{1})\Gamma(\alpha_{2})\Gamma(\alpha_{3})}{\Gamma(\alpha_{1}+\alpha_{2}+\alpha_{3})}.$
Using the copula representation given in Eq. (6) in the main paper, we have that
$f(x_{1},x_{2};\boldsymbol{\alpha})=f_{1}(x_{1};\boldsymbol{\alpha})\ f_{2}(x_{2};\boldsymbol{\alpha})\ c\left(u_{1},u_{2};\boldsymbol{\alpha}\right),$
where $u_{1}=F_{1}(x_{1})$ and $u_{2}=F_{2}(x_{2})$.
The marginals of a Dirichlet random vector are known to follow a Beta $X_{j}\sim\mathcal{B}eta\left(\alpha_{j},\sum_{i\neq j}\alpha_{i}\right)$, for $j=1,2$. Thus, using the copula representation we have that:
$c\left(u_{1},u_{2};\boldsymbol{\alpha}\right)$ $=\frac{f(x_{1},x_{2};\boldsymbol{\alpha})}{f_{1}(x_{1};\boldsymbol{\alpha})\ f_{2}(x_{2};\boldsymbol{\alpha})}$
[…]
$=\frac{(1-x_{1}-x_{2})^{\alpha_{3}-1}}{m\cdot(1-x_{1})^{\alpha_{2}+\alpha_{3}-1}(1-x_{2})^{\alpha_{1}+\alpha_{3}-1}}$
$=\frac{(1-F_{1}^{-1}(u_{1})-F_{2}^{-1}(u_{2}))^{\alpha_{3}-1}}{m\cdot(1-F_{1}^{-1}(u_{1}))^{\alpha_{2}+\alpha_{3}-1}(1-F_{2}^{-1}(u_{2}))^{\alpha_{1}+\alpha_{3}-1}},$ (8)
where $m=\frac{\Gamma(\alpha_{3})\Gamma(\alpha_{1}+\alpha_{2}+\alpha_{3})}{\Gamma(\alpha_{2}+\alpha_{3})\Gamma(\alpha_{1}+\alpha_{3})}$ denotes the normalizing constant and $F_{j}^{-1}$ is the inverse CDF of $X_{j}$.
Assume now that $\alpha_{1}=1,\alpha_{2}=1,\alpha_{3}=a$, with $a$ being a positive scalar. If $X_{j}$ is Beta distributed, then $F_{j}$ can be expressed in terms of the incomplete Beta function, denoted by $I_{j}$, which gives a closed form solution for both $u_{1}$ and $u_{2}$, and therefore $x_{1}$ and $x_{2}$:
$u_{j}=F_{j}(x_{j})=I_{j}(x_{j};\alpha_{1}=1,\alpha_{2}=1,\alpha_{3}=a)$ $=1-(1-x_{j})^{a+1},\quad j=1,2$
$x_{j}$ $=1-(1-u_{j})^{\frac{1}{a+1}},\quad j=1,2.$ (9)
Taken together, Eq. (A) and Eq. (A) give us, up to the normalizing constant $m$,
$c\left(u_{1},u_{2};\alpha_{1}=1,\alpha_{2}=1,\alpha_{3}=a\right)\propto\frac{\left[(1-u_{1})^{\frac{1}{a+1}}+(1-u_{2})^{\frac{1}{a+1}}-1\right]^{a-1}}{\left[(1-u_{1})(1-u_{2})\right]^{\frac{a}{a+1}}}.$
Since $x_{1},x_{2}\in[0,1]$, and $x_{1}+x_{2}\leq 1$, we also have the following constraints on $u_{1},u_{2}$, determined by Eq. (A):
$u_{1},u_{2}\in[0,1],\quad(1-u_{1})^{\frac{1}{a+1}}+(1-u_{2})^{\frac{1}{a+1}}\geq 1.$
Appendix B Supporting Details for Hyperparameter Choices
In this section, we illustrate the choices we made for the hyperparameters of the different measures evaluated in this paper, which we reiterate below. We support our choices with simulation studies in which we compare the performances of our measures for different values of the hyperparameters defined on an appropriate grid.
$M_{0}$:KDE
Direct estimation of the bivariate density. We use KDE, with Gaussian kernel and bandwidth selection based on the asympotically optimal solution proposed in Chacón et al., (2011), where its adequacy is shown in general settings, including Gaussian mixture models. Thus, no simulations are performed to empirically choose the hyperparameters.
$M^{\text{NPCop}}_{0}$:DE
Nonparametric indirect density estimation with copula. We use standard KDE with the same optimal bandwidth of Chacón et al., (2011) for the univariate marginals, and KDE with the transformation local likelihood estimator and nearest-neighbor bandwidth for the copula density (Nagler and Czado,, 2016). No simulations are performed to empirically choose the hyperparameters.
$M^{\text{PCop}}_{0}$:DE
Parametric indirect density estimation with copula. We adopt a fully parametric approach (with maximum likelihood fitting) to estimate both marginals and the copula. For the copula model, we select the best model using the AIC criterion; no misspecification is introduced for the marginals. No hyperpameters are involved; thus, no simulations are required.
$M_{1}$:$k$NN-Eucl
Cumulative Euclidean distances from the $k$NNs. We sum the Euclidean distances between each point and its $k$ neighbors defined according to the Euclidean metric. The choice of the hyperparameter $k$ is based on an extensive cross-validation procedure, leading to a general rule of thumb aligned with the existing literature: $k=\left[\sqrt{n/2}\right]$, with $\left[x\right]$ the integer closest to $x$. The supporting simulation results are given in Figures 9-12.
$M_{2}$:$k$NN-CDF
Cumulative CDF distances from the $k$NNs. We sum the CDF distances between each point and its $k$ neighbors, which are defined according to the Euclidean metric. The choice of the hyperparameter $k$ is based on an extensive cross-validation procedure, suggesting a uniform choice across different scenarios and sample sizes: $k=30$. The supporting simulation results are given in Figures 13-16.
$M_{3}$:$\epsilon$-CDF
This represents the $\epsilon$-neighborhood multivariate CDF distance, which uses empirical CDF as the CDF estimator. The optimal choice of the hyperparameter $\epsilon$ is based on an extensive cross-validation procedure, leading to the following heuristic for S1-S16 as a result of an exponential decay model fit with respect to the sample size: $\epsilon=\exp\left(2.13-0.3\log(n)\right)$. For S1-S16, the supporting simulation results are given in Figures 17-20. For S17 (Dirichlet scenario), performances are robust to sample size, with optimal $\epsilon=0.10$; see Figure 21.
$M^{\text{NPCop}}_{3}$:$\epsilon$-CDF
Fully nonparametric indirect estimation of the $\epsilon$-CDF measure with copula. We use the empirical CDF for estimating the univariate marginals, and KDE with the transformation local likelihood estimator and nearest-neighbor bandwidth for the copula density. The optimal choice of the hyperparameter $\epsilon$ follows the same strategy as the previous measure, leading to the following heuristic for S1-S16: $\epsilon=\exp\left(1.74-0.26\log(n)\right)$. For S1-S16, the supporting simulation results are given in Figures 22-25. For S17 (Dirichlet scenario), performances depend on the sample size, according to a heuristic given by: $\epsilon=\exp\left(-1.22-0.23\log(n)\right)$. This relationship is obtained by fitting a nonlinear regression (exponential decay model), with the empirical optimal $\epsilon$ and the sample size $n$ as dependent and independent variables, respectively; see Figure 26.
$M^{\text{PCop}}_{3}$:$\epsilon$-CDF
Parametric indirect CDF estimation with copula. We adopt a fully parametric approach (with maximum likelihood fitting) to estimate both marginals and the copula. For the copula model, we select the best model using the AIC criterion; no misspecification is introduced for the marginals. The optimal choice of the hyperparameter $\epsilon$ follows the heuristic $\epsilon=\exp\left(1.60-0.41\log(n)\right)$ for S1-S16 (see Figures 27-30). For S17 (Dirichlet scenario), performances are robust to sample size, with optimal $\epsilon=0.02$; see Figure 31.
Figure 9: Performance results in all evaluated scenarios for different values of the hyperparameter $k$ in $M_{1}$:$k$NN-Eucl. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=5$, with $n=50$.
Figure 10: Performance results in all evaluated scenarios for different values of the hyperparameter $k$ in $M_{1}$:$k$NN-Eucl. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=7$, with $n=100$.
Figure 11: Performance results in all evaluated scenarios for different values of the hyperparameter $k$ in $M_{1}$:$k$NN-Eucl. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=16$, with $n=500$.
Figure 12: Performance results in all evaluated scenarios for different values of the hyperparameter $k$ in $M_{1}$:$k$NN-Eucl. Gray vertical line refers to the rule of thumb for the optimal choice: $k=\left[\sqrt{n/2}\right]=22$, with $n=1000$.
Figure 13: Performance results in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $k$ in $M_{2}$:$k$NN-CDF. Gray vertical line refers to the final choice: $k=30$.
Figure 14: Performance results in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $k$ in $M_{2}$:$k$NN-CDF. Gray vertical line refers to the final choice: $k=30$.
Figure 15: Performance results in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $k$ in $M_{2}$:$k$NN-CDF. Gray vertical line refers to the final choice: $k=30$.
Figure 16: Performance results in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $k$ in $M_{2}$:$k$NN-CDF. Gray vertical line refers to the final choice: $k=30$.
Figure 17: Performance results in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $\epsilon$ in $M_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log(n)\right)=2.60$.
Figure 18: Performance results in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $\epsilon$ in $M_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log(n)\right)=2.11$.
Figure 19: Performance results in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $\epsilon$ in $M_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log(n)\right)=1.30$.
Figure 20: Performance results in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $\epsilon$ in $M_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(2.13-0.3\log(n)\right)=1.06$.
Figure 21: Performance results in the Dirichlet scenario for all sample sizes and for different values of the hyperparameter $\epsilon$ in $M_{3}$:$\epsilon$-CDF. Gray vertical line reflects the optimal choice given by $\epsilon=0.10$.
Figure 22: Performance results in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log(n)\right)\approx 2$.
Figure 23: Performance results in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log(n)\right)=1.72$.
Figure 24: Performance results in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log(n)\right)=1.13$.
Figure 25: Performance results in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log(n)\right)=0.95$.
Figure 26: Performance results in the Dirichlet scenario for all sample sizes and for different values of the hyperparameter $\epsilon$ in $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF. The optimal choice reflects the following heuristic: $\epsilon=\exp\left(-1.22-0.23\log(n)\right)$.
Figure 27: Performance results in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{PCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log(n)\right)\approx 1$.
Figure 28: Performance results in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{PCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log(n)\right)=0.75$.
Figure 29: Performance results in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{PCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log(n)\right)=0.39$.
Figure 30: Performance results in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{PCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log(n)\right)=0.29$.
Figure 31: Performance results in the Dirichlet scenario for all sample sizes and for different values of the hyperparameter $\epsilon$ in $M^{\text{PCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the optimal choice given by $\epsilon=0.02$.
Appendix C Simulation results for all scenarios
In this section, we report the full set of simulation results (for all scenarios and sample sizes) that were omitted in the main paper for the sake of space.
Table 2: Performance results in terms of mean (standard deviation) of the compared methods in scenario S1 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.035 (0.023) 0.029 (0.020) 0.031 (0.021) 0.034 (0.023) 0.057 (0.028) 0.030 (0.020) 0.034 (0.021) 0.030 (0.020)
FPR 0.024 (0.018) 0.020 (0.019) 0.021 (0.019) 0.025 (0.020) 0.035 (0.020) 0.021 (0.019) 0.023 (0.018) 0.021 (0.019)
FNR 0.175 (0.234) 0.120 (0.193) 0.146 (0.221) 0.133 (0.211) 0.452 (0.344) 0.126 (0.196) 0.158 (0.223) 0.122 (0.192)
Accuracy 0.965 (0.023) 0.971 (0.020) 0.969 (0.021) 0.966 (0.023) 0.943 (0.028) 0.970 (0.020) 0.966 (0.021) 0.970 (0.020)
F1 1.628 (0.263) 1.681 (0.263) 1.661 (0.272) 1.649 (0.267) 1.416 (0.305) 1.675 (0.262) 1.643 (0.260) 1.678 (0.261)
MCC 0.698 (0.197) 0.756 (0.176) 0.733 (0.199) 0.722 (0.192) 0.463 (0.298) 0.749 (0.178) 0.715 (0.188) 0.752 (0.174)
$n=100$
ERR 0.026 (0.015) 0.021 (0.013) 0.023 (0.014) 0.025 (0.016) 0.038 (0.017) 0.020 (0.014) 0.025 (0.014) 0.020 (0.013)
FPR 0.014 (0.012) 0.011 (0.012) 0.012 (0.013) 0.016 (0.013) 0.020 (0.014) 0.011 (0.012) 0.013 (0.012) 0.011 (0.012)
FNR 0.204 (0.195) 0.152 (0.165) 0.181 (0.179) 0.150 (0.174) 0.359 (0.224) 0.143 (0.165) 0.185 (0.180) 0.143 (0.162)
Accuracy 0.974 (0.015) 0.979 (0.013) 0.977 (0.014) 0.975 (0.016) 0.962 (0.017) 0.980 (0.014) 0.975 (0.014) 0.980 (0.013)
F1 1.707 (0.183) 1.762 (0.173) 1.737 (0.186) 1.728 (0.186) 1.573 (0.217) 1.769 (0.174) 1.726 (0.174) 1.770 (0.171)
MCC 0.733 (0.158) 0.788 (0.139) 0.762 (0.159) 0.754 (0.156) 0.593 (0.209) 0.796 (0.139) 0.752 (0.146) 0.796 (0.136)
$n=500$
ERR 0.015 (0.006) 0.013 (0.006) 0.018 (0.007) 0.017 (0.006) 0.015 (0.007) 0.009 (0.006) 0.014 (0.006) 0.009 (0.006)
FPR 0.008 (0.005) 0.007 (0.006) 0.009 (0.007) 0.010 (0.006) 0.008 (0.006) 0.004 (0.005) 0.007 (0.006) 0.005 (0.006)
FNR 0.135 (0.092) 0.117 (0.079) 0.173 (0.075) 0.132 (0.089) 0.139 (0.088) 0.073 (0.087) 0.125 (0.086) 0.075 (0.086)
Accuracy 0.985 (0.006) 0.987 (0.006) 0.982 (0.007) 0.983 (0.006) 0.985 (0.007) 0.991 (0.006) 0.986 (0.006) 0.991 (0.006)
F1 1.842 (0.061) 1.863 (0.063) 1.807 (0.080) 1.826 (0.066) 1.840 (0.075) 1.908 (0.061) 1.854 (0.062) 1.906 (0.061)
MCC 0.846 (0.058) 0.867 (0.059) 0.811 (0.077) 0.830 (0.063) 0.844 (0.072) 0.912 (0.055) 0.858 (0.058) 0.910 (0.055)
$n=1000$
ERR 0.012 (0.004) 0.011 (0.004) 0.018 (0.005) 0.014 (0.004) 0.009 (0.004) 0.006 (0.004) 0.011 (0.004) 0.006 (0.004)
FPR 0.006 (0.004) 0.006 (0.004) 0.010 (0.005) 0.008 (0.004) 0.005 (0.004) 0.003 (0.004) 0.006 (0.004) 0.003 (0.004)
FNR 0.113 (0.067) 0.100 (0.056) 0.181 (0.053) 0.115 (0.066) 0.086 (0.065) 0.052 (0.064) 0.098 (0.060) 0.053 (0.064)
Accuracy 0.988 (0.004) 0.989 (0.004) 0.982 (0.005) 0.986 (0.004) 0.991 (0.004) 0.994 (0.004) 0.989 (0.004) 0.994 (0.004)
F1 1.873 (0.048) 1.887 (0.043) 1.803 (0.060) 1.855 (0.045) 1.901 (0.045) 1.937 (0.042) 1.888 (0.040) 1.935 (0.042)
MCC 0.875 (0.046) 0.889 (0.040) 0.806 (0.059) 0.858 (0.044) 0.903 (0.043) 0.939 (0.039) 0.891 (0.038) 0.938 (0.039)
Table 3: Performance results in terms of mean (standard deviation) of the compared methods in scenario S2 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.043 (0.025) 0.043 (0.024) 0.055 (0.026) 0.049 (0.026) 0.042 (0.024) 0.040 (0.023) 0.044 (0.024) 0.037 (0.022)
FPR 0.028 (0.018) 0.028 (0.018) 0.034 (0.019) 0.033 (0.022) 0.027 (0.019) 0.026 (0.019) 0.029 (0.019) 0.025 (0.019)
FNR 0.252 (0.280) 0.251 (0.279) 0.402 (0.330) 0.292 (0.297) 0.247 (0.276) 0.219 (0.262) 0.274 (0.281) 0.195 (0.243)
Accuracy 0.957 (0.025) 0.957 (0.024) 0.945 (0.026) 0.951 (0.026) 0.958 (0.024) 0.960 (0.023) 0.956 (0.024) 0.963 (0.022)
F1 1.549 (0.274) 1.553 (0.275) 1.435 (0.286) 1.509 (0.284) 1.557 (0.276) 1.579 (0.273) 1.535 (0.274) 1.600 (0.269)
MCC 0.623 (0.224) 0.627 (0.224) 0.495 (0.270) 0.580 (0.246) 0.632 (0.224) 0.657 (0.213) 0.607 (0.228) 0.680 (0.198)
$n=100$
ERR 0.031 (0.017) 0.032 (0.016) 0.048 (0.019) 0.036 (0.016) 0.029 (0.016) 0.026 (0.016) 0.032 (0.016) 0.025 (0.015)
FPR 0.017 (0.012) 0.017 (0.013) 0.025 (0.013) 0.020 (0.014) 0.015 (0.012) 0.014 (0.013) 0.017 (0.013) 0.013 (0.012)
FNR 0.253 (0.212) 0.263 (0.206) 0.450 (0.235) 0.282 (0.201) 0.228 (0.203) 0.198 (0.193) 0.265 (0.203) 0.186 (0.188)
Accuracy 0.969 (0.017) 0.968 (0.016) 0.952 (0.019) 0.964 (0.016) 0.971 (0.016) 0.974 (0.016) 0.968 (0.016) 0.975 (0.015)
F1 1.655 (0.190) 1.649 (0.191) 1.478 (0.214) 1.615 (0.193) 1.682 (0.191) 1.712 (0.187) 1.649 (0.190) 1.723 (0.185)
MCC 0.680 (0.172) 0.673 (0.173) 0.497 (0.210) 0.640 (0.176) 0.707 (0.169) 0.738 (0.162) 0.673 (0.172) 0.749 (0.157)
$n=500$
ERR 0.016 (0.006) 0.018 (0.006) 0.050 (0.009) 0.020 (0.007) 0.014 (0.006) 0.011 (0.006) 0.017 (0.006) 0.011 (0.006)
FPR 0.008 (0.005) 0.010 (0.006) 0.026 (0.006) 0.011 (0.006) 0.007 (0.005) 0.005 (0.006) 0.009 (0.005) 0.005 (0.006)
FNR 0.144 (0.092) 0.172 (0.092) 0.498 (0.094) 0.179 (0.091) 0.123 (0.090) 0.093 (0.088) 0.157 (0.091) 0.092 (0.088)
Accuracy 0.984 (0.006) 0.982 (0.006) 0.950 (0.009) 0.980 (0.007) 0.986 (0.006) 0.989 (0.006) 0.983 (0.006) 0.989 (0.006)
F1 1.834 (0.065) 1.805 (0.069) 1.469 (0.098) 1.789 (0.070) 1.856 (0.062) 1.887 (0.062) 1.821 (0.066) 1.888 (0.063)
MCC 0.838 (0.061) 0.810 (0.067) 0.472 (0.098) 0.794 (0.068) 0.860 (0.058) 0.892 (0.057) 0.825 (0.063) 0.893 (0.058)
$n=1000$
ERR 0.011 (0.004) 0.015 (0.004) 0.051 (0.007) 0.016 (0.004) 0.010 (0.004) 0.007 (0.004) 0.013 (0.004) 0.007 (0.004)
FPR 0.006 (0.004) 0.008 (0.004) 0.027 (0.004) 0.009 (0.004) 0.005 (0.004) 0.004 (0.004) 0.007 (0.004) 0.004 (0.004)
FNR 0.105 (0.067) 0.137 (0.069) 0.505 (0.066) 0.139 (0.067) 0.090 (0.066) 0.065 (0.066) 0.120 (0.068) 0.065 (0.066)
Accuracy 0.989 (0.004) 0.985 (0.004) 0.949 (0.007) 0.984 (0.004) 0.990 (0.004) 0.993 (0.004) 0.987 (0.004) 0.993 (0.004)
F1 1.879 (0.044) 1.846 (0.046) 1.465 (0.070) 1.836 (0.047) 1.896 (0.042) 1.922 (0.042) 1.864 (0.045) 1.922 (0.042)
MCC 0.882 (0.042) 0.848 (0.045) 0.466 (0.069) 0.838 (0.046) 0.898 (0.040) 0.924 (0.039) 0.867 (0.043) 0.924 (0.039)
Table 4: Performance results in terms of mean (standard deviation) of the compared methods in scenario S3 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.046 (0.027) 0.045 (0.025) 0.062 (0.028) 0.051 (0.026) 0.043 (0.024) 0.043 (0.025) 0.044 (0.024) 0.040 (0.023)
FPR 0.030 (0.019) 0.029 (0.018) 0.038 (0.018) 0.035 (0.022) 0.028 (0.019) 0.028 (0.019) 0.029 (0.019) 0.026 (0.019)
FNR 0.289 (0.304) 0.281 (0.294) 0.492 (0.340) 0.299 (0.303) 0.266 (0.283) 0.261 (0.287) 0.270 (0.288) 0.229 (0.267)
Accuracy 0.954 (0.027) 0.955 (0.025) 0.938 (0.028) 0.949 (0.026) 0.957 (0.024) 0.957 (0.025) 0.956 (0.024) 0.960 (0.023)
F1 1.522 (0.287) 1.529 (0.278) 1.363 (0.284) 1.493 (0.280) 1.545 (0.279) 1.549 (0.285) 1.540 (0.279) 1.575 (0.278)
MCC 0.593 (0.248) 0.601 (0.235) 0.415 (0.279) 0.565 (0.243) 0.618 (0.232) 0.623 (0.238) 0.613 (0.233) 0.651 (0.221)
$n=100$
ERR 0.033 (0.017) 0.033 (0.017) 0.053 (0.019) 0.036 (0.017) 0.029 (0.016) 0.029 (0.016) 0.031 (0.016) 0.027 (0.016)
FPR 0.018 (0.012) 0.018 (0.013) 0.028 (0.013) 0.021 (0.014) 0.016 (0.012) 0.015 (0.013) 0.017 (0.013) 0.015 (0.013)
FNR 0.271 (0.215) 0.273 (0.206) 0.515 (0.239) 0.277 (0.206) 0.232 (0.202) 0.227 (0.198) 0.255 (0.205) 0.215 (0.195)
Accuracy 0.967 (0.017) 0.967 (0.017) 0.947 (0.019) 0.964 (0.017) 0.971 (0.016) 0.971 (0.016) 0.969 (0.016) 0.973 (0.016)
F1 1.636 (0.194) 1.637 (0.194) 1.416 (0.218) 1.612 (0.199) 1.674 (0.189) 1.682 (0.191) 1.656 (0.195) 1.693 (0.192)
MCC 0.661 (0.176) 0.662 (0.175) 0.434 (0.216) 0.638 (0.180) 0.701 (0.165) 0.708 (0.167) 0.681 (0.175) 0.720 (0.166)
$n=500$
ERR 0.017 (0.006) 0.019 (0.006) 0.056 (0.009) 0.020 (0.007) 0.014 (0.006) 0.012 (0.006) 0.016 (0.006) 0.012 (0.006)
FPR 0.009 (0.005) 0.010 (0.005) 0.029 (0.006) 0.011 (0.006) 0.007 (0.005) 0.006 (0.005) 0.008 (0.005) 0.006 (0.005)
FNR 0.155 (0.094) 0.173 (0.091) 0.558 (0.093) 0.171 (0.089) 0.124 (0.090) 0.102 (0.090) 0.143 (0.092) 0.102 (0.090)
Accuracy 0.983 (0.006) 0.981 (0.006) 0.944 (0.009) 0.980 (0.007) 0.986 (0.006) 0.988 (0.006) 0.984 (0.006) 0.988 (0.006)
F1 1.822 (0.064) 1.804 (0.067) 1.408 (0.097) 1.794 (0.068) 1.855 (0.061) 1.878 (0.063) 1.835 (0.063) 1.878 (0.062)
MCC 0.826 (0.062) 0.808 (0.065) 0.410 (0.097) 0.799 (0.066) 0.859 (0.058) 0.882 (0.059) 0.839 (0.061) 0.883 (0.058)
$n=1000$
ERR 0.012 (0.004) 0.015 (0.004) 0.056 (0.007) 0.016 (0.004) 0.010 (0.004) 0.008 (0.004) 0.012 (0.004) 0.008 (0.004)
FPR 0.007 (0.004) 0.008 (0.004) 0.030 (0.004) 0.009 (0.004) 0.005 (0.004) 0.004 (0.004) 0.006 (0.004) 0.004 (0.004)
FNR 0.117 (0.068) 0.141 (0.068) 0.561 (0.065) 0.141 (0.067) 0.092 (0.065) 0.071 (0.065) 0.111 (0.067) 0.071 (0.065)
Accuracy 0.988 (0.004) 0.985 (0.004) 0.944 (0.007) 0.984 (0.004) 0.990 (0.004) 0.992 (0.004) 0.988 (0.004) 0.992 (0.004)
F1 1.869 (0.043) 1.843 (0.045) 1.407 (0.069) 1.834 (0.046) 1.895 (0.040) 1.916 (0.041) 1.874 (0.042) 1.917 (0.041)
MCC 0.871 (0.041) 0.846 (0.044) 0.408 (0.069) 0.836 (0.045) 0.897 (0.038) 0.919 (0.038) 0.876 (0.040) 0.919 (0.038)
Table 5: Performance results in terms of mean (standard deviation) of the compared methods in scenario S4 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.061 (0.030) 0.060 (0.028) 0.081 (0.030) 0.068 (0.032) 0.050 (0.027) 0.050 (0.027) 0.055 (0.027) 0.049 (0.027)
FPR 0.038 (0.018) 0.037 (0.018) 0.048 (0.016) 0.046 (0.024) 0.031 (0.019) 0.032 (0.019) 0.034 (0.019) 0.031 (0.019)
FNR 0.464 (0.342) 0.454 (0.336) 0.709 (0.321) 0.444 (0.347) 0.335 (0.314) 0.337 (0.314) 0.400 (0.324) 0.326 (0.310)
Accuracy 0.939 (0.030) 0.940 (0.028) 0.919 (0.030) 0.932 (0.032) 0.950 (0.027) 0.950 (0.027) 0.945 (0.027) 0.951 (0.027)
F1 1.379 (0.287) 1.391 (0.287) 1.188 (0.260) 1.363 (0.284) 1.485 (0.292) 1.481 (0.290) 1.434 (0.287) 1.490 (0.289)
MCC 0.438 (0.277) 0.451 (0.275) 0.221 (0.273) 0.426 (0.274) 0.559 (0.257) 0.555 (0.255) 0.500 (0.265) 0.564 (0.252)
$n=100$
ERR 0.046 (0.020) 0.043 (0.020) 0.069 (0.021) 0.050 (0.021) 0.035 (0.017) 0.034 (0.017) 0.038 (0.018) 0.034 (0.017)
FPR 0.024 (0.012) 0.023 (0.013) 0.036 (0.011) 0.030 (0.015) 0.019 (0.013) 0.018 (0.013) 0.020 (0.013) 0.018 (0.012)
FNR 0.410 (0.242) 0.388 (0.238) 0.683 (0.233) 0.380 (0.237) 0.297 (0.218) 0.290 (0.218) 0.335 (0.224) 0.280 (0.217)
Accuracy 0.954 (0.020) 0.957 (0.020) 0.931 (0.021) 0.950 (0.021) 0.965 (0.017) 0.966 (0.017) 0.962 (0.018) 0.966 (0.017)
F1 1.506 (0.207) 1.529 (0.209) 1.259 (0.209) 1.497 (0.203) 1.617 (0.196) 1.623 (0.196) 1.581 (0.200) 1.631 (0.194)
MCC 0.526 (0.202) 0.550 (0.203) 0.270 (0.213) 0.520 (0.195) 0.640 (0.182) 0.647 (0.182) 0.603 (0.189) 0.655 (0.179)
$n=500$
ERR 0.027 (0.007) 0.025 (0.007) 0.069 (0.009) 0.030 (0.008) 0.019 (0.006) 0.014 (0.006) 0.019 (0.007) 0.014 (0.006)
FPR 0.014 (0.006) 0.013 (0.006) 0.036 (0.005) 0.018 (0.006) 0.010 (0.006) 0.008 (0.006) 0.010 (0.006) 0.008 (0.006)
FNR 0.253 (0.095) 0.234 (0.094) 0.686 (0.084) 0.256 (0.099) 0.177 (0.087) 0.131 (0.089) 0.180 (0.091) 0.131 (0.089)
Accuracy 0.973 (0.007) 0.975 (0.007) 0.931 (0.009) 0.970 (0.008) 0.981 (0.006) 0.986 (0.006) 0.981 (0.007) 0.986 (0.006)
F1 1.719 (0.079) 1.740 (0.077) 1.274 (0.087) 1.691 (0.080) 1.800 (0.072) 1.846 (0.069) 1.796 (0.072) 1.847 (0.069)
MCC 0.723 (0.078) 0.744 (0.075) 0.275 (0.087) 0.696 (0.079) 0.804 (0.070) 0.851 (0.065) 0.800 (0.069) 0.851 (0.065)
$n=1000$
ERR 0.022 (0.005) 0.020 (0.005) 0.069 (0.007) 0.025 (0.005) 0.015 (0.004) 0.010 (0.004) 0.014 (0.004) 0.010 (0.004)
FPR 0.011 (0.004) 0.010 (0.004) 0.036 (0.003) 0.015 (0.004) 0.008 (0.004) 0.005 (0.004) 0.008 (0.004) 0.005 (0.004)
FNR 0.211 (0.067) 0.190 (0.068) 0.684 (0.060) 0.225 (0.072) 0.140 (0.059) 0.092 (0.065) 0.136 (0.066) 0.092 (0.065)
Accuracy 0.978 (0.005) 0.980 (0.005) 0.931 (0.007) 0.975 (0.005) 0.985 (0.004) 0.990 (0.004) 0.986 (0.004) 0.990 (0.004)
F1 1.771 (0.052) 1.792 (0.051) 1.277 (0.062) 1.737 (0.052) 1.845 (0.048) 1.894 (0.042) 1.848 (0.044) 1.894 (0.042)
MCC 0.773 (0.052) 0.794 (0.050) 0.278 (0.062) 0.740 (0.051) 0.847 (0.046) 0.896 (0.040) 0.850 (0.043) 0.897 (0.040)
Table 6: Performance results in terms of mean (standard deviation) of the compared methods in scenario S5 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.033 (0.021) 0.030 (0.020) 0.034 (0.022) 0.034 (0.022) 0.054 (0.026) 0.031 (0.020) 0.034 (0.021) 0.030 (0.020)
FPR 0.023 (0.019) 0.021 (0.020) 0.023 (0.020) 0.025 (0.020) 0.034 (0.020) 0.021 (0.020) 0.023 (0.019) 0.021 (0.020)
FNR 0.150 (0.210) 0.128 (0.192) 0.179 (0.238) 0.132 (0.206) 0.441 (0.332) 0.132 (0.193) 0.166 (0.221) 0.125 (0.189)
Accuracy 0.967 (0.021) 0.970 (0.020) 0.966 (0.022) 0.966 (0.022) 0.946 (0.026) 0.969 (0.020) 0.966 (0.021) 0.970 (0.020)
F1 1.639 (0.274) 1.663 (0.279) 1.625 (0.291) 1.638 (0.283) 1.426 (0.312) 1.659 (0.277) 1.626 (0.274) 1.666 (0.278)
MCC 0.728 (0.177) 0.754 (0.174) 0.711 (0.212) 0.727 (0.188) 0.485 (0.298) 0.750 (0.172) 0.713 (0.184) 0.757 (0.171)
$n=100$
ERR 0.025 (0.015) 0.021 (0.013) 0.026 (0.014) 0.024 (0.015) 0.037 (0.017) 0.020 (0.013) 0.024 (0.014) 0.020 (0.013)
FPR 0.013 (0.012) 0.011 (0.012) 0.014 (0.013) 0.016 (0.013) 0.020 (0.013) 0.011 (0.012) 0.013 (0.012) 0.011 (0.012)
FNR 0.184 (0.192) 0.148 (0.164) 0.212 (0.187) 0.137 (0.163) 0.338 (0.221) 0.141 (0.166) 0.176 (0.179) 0.137 (0.162)
Accuracy 0.975 (0.015) 0.979 (0.013) 0.974 (0.014) 0.976 (0.015) 0.963 (0.017) 0.980 (0.013) 0.976 (0.014) 0.980 (0.013)
F1 1.726 (0.174) 1.764 (0.166) 1.708 (0.186) 1.741 (0.175) 1.590 (0.211) 1.771 (0.165) 1.735 (0.166) 1.774 (0.164)
MCC 0.749 (0.156) 0.787 (0.140) 0.729 (0.170) 0.764 (0.151) 0.608 (0.205) 0.794 (0.139) 0.758 (0.145) 0.798 (0.137)
$n=500$
ERR 0.014 (0.006) 0.012 (0.005) 0.024 (0.007) 0.016 (0.006) 0.015 (0.006) 0.009 (0.006) 0.013 (0.006) 0.009 (0.006)
FPR 0.007 (0.005) 0.006 (0.006) 0.013 (0.007) 0.010 (0.006) 0.008 (0.006) 0.005 (0.006) 0.007 (0.005) 0.005 (0.006)
FNR 0.123 (0.095) 0.109 (0.082) 0.238 (0.080) 0.118 (0.087) 0.135 (0.087) 0.076 (0.090) 0.113 (0.091) 0.075 (0.089)
Accuracy 0.986 (0.006) 0.988 (0.005) 0.976 (0.007) 0.984 (0.006) 0.985 (0.006) 0.991 (0.006) 0.987 (0.006) 0.991 (0.006)
F1 1.853 (0.061) 1.869 (0.062) 1.739 (0.091) 1.837 (0.061) 1.842 (0.067) 1.903 (0.060) 1.864 (0.059) 1.903 (0.060)
MCC 0.858 (0.057) 0.874 (0.057) 0.743 (0.089) 0.842 (0.057) 0.847 (0.064) 0.907 (0.054) 0.869 (0.055) 0.908 (0.054)
$n=1000$
ERR 0.010 (0.004) 0.010 (0.004) 0.026 (0.006) 0.013 (0.004) 0.010 (0.004) 0.006 (0.004) 0.009 (0.004) 0.006 (0.004)
FPR 0.005 (0.004) 0.005 (0.004) 0.014 (0.005) 0.008 (0.004) 0.005 (0.004) 0.003 (0.004) 0.005 (0.004) 0.003 (0.004)
FNR 0.092 (0.067) 0.091 (0.057) 0.257 (0.059) 0.105 (0.067) 0.087 (0.064) 0.053 (0.064) 0.084 (0.065) 0.053 (0.063)
Accuracy 0.990 (0.004) 0.990 (0.004) 0.974 (0.006) 0.987 (0.004) 0.990 (0.004) 0.994 (0.004) 0.991 (0.004) 0.994 (0.004)
F1 1.893 (0.042) 1.894 (0.041) 1.723 (0.068) 1.864 (0.041) 1.898 (0.040) 1.933 (0.039) 1.901 (0.038) 1.934 (0.039)
MCC 0.896 (0.041) 0.897 (0.040) 0.725 (0.067) 0.866 (0.040) 0.900 (0.038) 0.936 (0.036) 0.903 (0.036) 0.936 (0.037)
Table 7: Performance results in terms of mean (standard deviation) of the compared methods in scenario S6 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.042 (0.025) 0.042 (0.023) 0.062 (0.027) 0.049 (0.025) 0.042 (0.024) 0.039 (0.023) 0.043 (0.023) 0.036 (0.022)
FPR 0.027 (0.019) 0.028 (0.018) 0.038 (0.018) 0.033 (0.021) 0.027 (0.019) 0.026 (0.019) 0.028 (0.019) 0.025 (0.019)
FNR 0.242 (0.287) 0.248 (0.270) 0.495 (0.338) 0.296 (0.293) 0.243 (0.276) 0.212 (0.255) 0.258 (0.274) 0.186 (0.238)
Accuracy 0.958 (0.025) 0.958 (0.023) 0.938 (0.027) 0.951 (0.025) 0.958 (0.024) 0.961 (0.023) 0.957 (0.023) 0.964 (0.022)
F1 1.560 (0.287) 1.557 (0.274) 1.364 (0.286) 1.502 (0.279) 1.560 (0.280) 1.588 (0.277) 1.548 (0.276) 1.609 (0.275)
MCC 0.636 (0.236) 0.632 (0.219) 0.415 (0.281) 0.574 (0.240) 0.636 (0.227) 0.666 (0.212) 0.623 (0.224) 0.690 (0.201)
$n=100$
ERR 0.030 (0.016) 0.031 (0.016) 0.058 (0.020) 0.035 (0.017) 0.029 (0.015) 0.026 (0.015) 0.032 (0.016) 0.025 (0.015)
FPR 0.016 (0.012) 0.017 (0.012) 0.030 (0.012) 0.020 (0.014) 0.015 (0.012) 0.014 (0.012) 0.017 (0.012) 0.013 (0.012)
FNR 0.244 (0.204) 0.259 (0.205) 0.559 (0.233) 0.278 (0.208) 0.229 (0.197) 0.206 (0.189) 0.262 (0.206) 0.191 (0.184)
Accuracy 0.970 (0.016) 0.969 (0.016) 0.943 (0.020) 0.965 (0.017) 0.971 (0.015) 0.974 (0.015) 0.968 (0.016) 0.975 (0.015)
F1 1.672 (0.185) 1.656 (0.185) 1.377 (0.211) 1.622 (0.194) 1.683 (0.175) 1.709 (0.180) 1.655 (0.186) 1.723 (0.176)
MCC 0.696 (0.165) 0.680 (0.167) 0.392 (0.213) 0.646 (0.178) 0.708 (0.153) 0.735 (0.155) 0.678 (0.169) 0.749 (0.149)
$n=500$
ERR 0.015 (0.006) 0.017 (0.006) 0.060 (0.009) 0.019 (0.006) 0.014 (0.006) 0.011 (0.005) 0.017 (0.006) 0.011 (0.005)
FPR 0.008 (0.005) 0.009 (0.005) 0.032 (0.006) 0.010 (0.006) 0.007 (0.005) 0.006 (0.005) 0.009 (0.005) 0.006 (0.005)
FNR 0.140 (0.089) 0.161 (0.094) 0.601 (0.090) 0.165 (0.092) 0.128 (0.090) 0.096 (0.085) 0.157 (0.093) 0.095 (0.085)
Accuracy 0.985 (0.006) 0.983 (0.006) 0.940 (0.009) 0.981 (0.006) 0.986 (0.006) 0.989 (0.005) 0.983 (0.006) 0.989 (0.005)
F1 1.838 (0.062) 1.816 (0.065) 1.363 (0.096) 1.804 (0.068) 1.851 (0.059) 1.884 (0.058) 1.820 (0.061) 1.886 (0.058)
MCC 0.843 (0.059) 0.820 (0.062) 0.365 (0.096) 0.808 (0.066) 0.855 (0.055) 0.888 (0.053) 0.825 (0.059) 0.890 (0.053)
$n=1000$
ERR 0.012 (0.004) 0.014 (0.004) 0.061 (0.007) 0.014 (0.004) 0.011 (0.004) 0.008 (0.004) 0.013 (0.004) 0.008 (0.004)
FPR 0.006 (0.004) 0.007 (0.004) 0.032 (0.004) 0.008 (0.004) 0.006 (0.004) 0.004 (0.004) 0.007 (0.004) 0.004 (0.004)
FNR 0.109 (0.066) 0.130 (0.066) 0.608 (0.066) 0.129 (0.065) 0.100 (0.065) 0.067 (0.063) 0.123 (0.066) 0.067 (0.063)
Accuracy 0.988 (0.004) 0.986 (0.004) 0.939 (0.007) 0.986 (0.004) 0.989 (0.004) 0.992 (0.004) 0.987 (0.004) 0.992 (0.004)
F1 1.876 (0.042) 1.854 (0.044) 1.357 (0.071) 1.849 (0.044) 1.886 (0.040) 1.920 (0.038) 1.861 (0.040) 1.920 (0.038)
MCC 0.878 (0.041) 0.856 (0.043) 0.358 (0.071) 0.851 (0.042) 0.888 (0.038) 0.922 (0.035) 0.863 (0.039) 0.923 (0.035)
Table 8: Performance results in terms of mean (standard deviation) of the compared methods in scenario S7 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.046 (0.026) 0.043 (0.025) 0.073 (0.030) 0.050 (0.027) 0.043 (0.025) 0.042 (0.024) 0.043 (0.025) 0.040 (0.024)
FPR 0.030 (0.019) 0.028 (0.018) 0.044 (0.017) 0.034 (0.023) 0.028 (0.019) 0.027 (0.019) 0.028 (0.019) 0.026 (0.019)
FNR 0.296 (0.301) 0.258 (0.279) 0.620 (0.334) 0.287 (0.294) 0.257 (0.286) 0.259 (0.277) 0.255 (0.277) 0.226 (0.265)
Accuracy 0.954 (0.026) 0.957 (0.025) 0.927 (0.030) 0.950 (0.027) 0.958 (0.025) 0.958 (0.024) 0.957 (0.025) 0.960 (0.024)
F1 1.522 (0.283) 1.551 (0.274) 1.267 (0.279) 1.511 (0.280) 1.555 (0.283) 1.555 (0.276) 1.556 (0.276) 1.582 (0.277)
MCC 0.589 (0.247) 0.621 (0.226) 0.303 (0.286) 0.581 (0.242) 0.625 (0.237) 0.625 (0.229) 0.626 (0.228) 0.655 (0.221)
$n=100$
ERR 0.034 (0.017) 0.033 (0.017) 0.065 (0.021) 0.036 (0.018) 0.030 (0.016) 0.029 (0.016) 0.031 (0.016) 0.028 (0.016)
FPR 0.018 (0.013) 0.017 (0.012) 0.034 (0.012) 0.021 (0.014) 0.016 (0.012) 0.015 (0.012) 0.016 (0.013) 0.014 (0.012)
FNR 0.286 (0.212) 0.274 (0.217) 0.636 (0.236) 0.272 (0.210) 0.245 (0.199) 0.234 (0.201) 0.257 (0.205) 0.222 (0.199)
Accuracy 0.966 (0.017) 0.967 (0.017) 0.935 (0.021) 0.964 (0.018) 0.970 (0.016) 0.971 (0.016) 0.969 (0.016) 0.972 (0.016)
F1 1.633 (0.195) 1.643 (0.194) 1.302 (0.211) 1.621 (0.196) 1.669 (0.180) 1.684 (0.190) 1.662 (0.191) 1.695 (0.189)
MCC 0.657 (0.180) 0.666 (0.178) 0.315 (0.214) 0.646 (0.179) 0.694 (0.159) 0.708 (0.169) 0.686 (0.173) 0.720 (0.166)
$n=500$
ERR 0.018 (0.006) 0.018 (0.006) 0.068 (0.009) 0.019 (0.006) 0.015 (0.006) 0.012 (0.005) 0.017 (0.006) 0.012 (0.005)
FPR 0.010 (0.005) 0.010 (0.005) 0.036 (0.005) 0.011 (0.006) 0.008 (0.005) 0.006 (0.005) 0.009 (0.005) 0.006 (0.006)
FNR 0.171 (0.091) 0.172 (0.091) 0.680 (0.088) 0.166 (0.090) 0.139 (0.090) 0.104 (0.087) 0.154 (0.090) 0.103 (0.087)
Accuracy 0.982 (0.006) 0.982 (0.006) 0.932 (0.009) 0.981 (0.006) 0.985 (0.006) 0.988 (0.005) 0.983 (0.006) 0.988 (0.005)
F1 1.806 (0.067) 1.806 (0.067) 1.282 (0.093) 1.798 (0.068) 1.839 (0.062) 1.876 (0.060) 1.824 (0.062) 1.877 (0.061)
MCC 0.811 (0.064) 0.810 (0.064) 0.283 (0.093) 0.802 (0.066) 0.844 (0.059) 0.881 (0.056) 0.828 (0.059) 0.882 (0.056)
$n=1000$
ERR 0.015 (0.004) 0.015 (0.004) 0.069 (0.007) 0.016 (0.004) 0.011 (0.004) 0.008 (0.004) 0.013 (0.004) 0.008 (0.004)
FPR 0.008 (0.004) 0.008 (0.004) 0.036 (0.004) 0.009 (0.004) 0.006 (0.004) 0.004 (0.004) 0.007 (0.004) 0.004 (0.004)
FNR 0.142 (0.067) 0.142 (0.067) 0.684 (0.063) 0.142 (0.065) 0.108 (0.067) 0.077 (0.066) 0.124 (0.067) 0.077 (0.066)
Accuracy 0.985 (0.004) 0.985 (0.004) 0.931 (0.007) 0.984 (0.004) 0.989 (0.004) 0.992 (0.004) 0.987 (0.004) 0.992 (0.004)
F1 1.845 (0.044) 1.845 (0.044) 1.279 (0.067) 1.833 (0.044) 1.880 (0.040) 1.913 (0.040) 1.864 (0.040) 1.913 (0.040)
MCC 0.847 (0.043) 0.847 (0.043) 0.280 (0.067) 0.835 (0.043) 0.882 (0.039) 0.915 (0.038) 0.866 (0.038) 0.916 (0.038)
Table 9: Performance results in terms of mean (standard deviation) of the compared methods in scenario S8 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.058 (0.029) 0.057 (0.027) 0.087 (0.031) 0.066 (0.031) 0.046 (0.025) 0.048 (0.025) 0.052 (0.027) 0.048 (0.025)
FPR 0.036 (0.018) 0.035 (0.018) 0.051 (0.015) 0.045 (0.025) 0.029 (0.019) 0.031 (0.018) 0.033 (0.019) 0.030 (0.018)
FNR 0.425 (0.339) 0.428 (0.335) 0.770 (0.295) 0.430 (0.344) 0.291 (0.299) 0.321 (0.306) 0.372 (0.327) 0.311 (0.299)
Accuracy 0.942 (0.029) 0.943 (0.027) 0.913 (0.031) 0.934 (0.031) 0.954 (0.025) 0.952 (0.025) 0.948 (0.027) 0.952 (0.025)
F1 1.413 (0.290) 1.414 (0.288) 1.140 (0.241) 1.378 (0.285) 1.522 (0.284) 1.497 (0.281) 1.460 (0.294) 1.504 (0.278)
MCC 0.471 (0.276) 0.473 (0.274) 0.163 (0.254) 0.439 (0.274) 0.594 (0.243) 0.566 (0.246) 0.524 (0.271) 0.574 (0.241)
$n=100$
ERR 0.042 (0.018) 0.041 (0.018) 0.076 (0.021) 0.047 (0.020) 0.033 (0.016) 0.033 (0.017) 0.036 (0.017) 0.033 (0.017)
FPR 0.022 (0.012) 0.021 (0.013) 0.040 (0.011) 0.028 (0.016) 0.017 (0.012) 0.017 (0.013) 0.019 (0.013) 0.017 (0.013)
FNR 0.377 (0.232) 0.369 (0.225) 0.760 (0.215) 0.360 (0.231) 0.275 (0.208) 0.281 (0.214) 0.316 (0.215) 0.278 (0.209)
Accuracy 0.958 (0.018) 0.959 (0.018) 0.924 (0.021) 0.953 (0.020) 0.967 (0.016) 0.967 (0.017) 0.964 (0.017) 0.967 (0.017)
F1 1.544 (0.203) 1.555 (0.202) 1.187 (0.194) 1.524 (0.206) 1.641 (0.188) 1.639 (0.199) 1.605 (0.199) 1.641 (0.197)
MCC 0.565 (0.195) 0.576 (0.194) 0.196 (0.200) 0.547 (0.196) 0.665 (0.171) 0.662 (0.184) 0.627 (0.186) 0.664 (0.180)
$n=500$
ERR 0.024 (0.007) 0.023 (0.007) 0.076 (0.010) 0.028 (0.008) 0.017 (0.006) 0.014 (0.006) 0.019 (0.006) 0.014 (0.006)
FPR 0.013 (0.006) 0.012 (0.006) 0.040 (0.005) 0.016 (0.006) 0.009 (0.006) 0.007 (0.006) 0.010 (0.006) 0.007 (0.006)
FNR 0.233 (0.092) 0.222 (0.091) 0.758 (0.082) 0.241 (0.097) 0.162 (0.089) 0.127 (0.088) 0.183 (0.089) 0.126 (0.088)
Accuracy 0.976 (0.007) 0.977 (0.007) 0.924 (0.010) 0.972 (0.008) 0.983 (0.006) 0.986 (0.006) 0.981 (0.006) 0.986 (0.006)
F1 1.744 (0.075) 1.755 (0.074) 1.200 (0.085) 1.711 (0.078) 1.818 (0.068) 1.854 (0.063) 1.796 (0.070) 1.855 (0.063)
MCC 0.748 (0.073) 0.759 (0.072) 0.201 (0.085) 0.716 (0.076) 0.822 (0.065) 0.858 (0.060) 0.800 (0.067) 0.860 (0.060)
$n=1000$
ERR 0.020 (0.005) 0.019 (0.005) 0.076 (0.007) 0.024 (0.005) 0.013 (0.004) 0.010 (0.004) 0.015 (0.004) 0.010 (0.004)
FPR 0.011 (0.004) 0.010 (0.004) 0.040 (0.003) 0.013 (0.004) 0.007 (0.004) 0.005 (0.004) 0.008 (0.004) 0.005 (0.004)
FNR 0.197 (0.068) 0.180 (0.067) 0.762 (0.059) 0.206 (0.070) 0.126 (0.064) 0.090 (0.064) 0.146 (0.062) 0.090 (0.064)
Accuracy 0.980 (0.005) 0.981 (0.005) 0.924 (0.007) 0.976 (0.005) 0.987 (0.004) 0.990 (0.004) 0.985 (0.004) 0.990 (0.004)
F1 1.787 (0.052) 1.804 (0.051) 1.196 (0.062) 1.758 (0.051) 1.861 (0.043) 1.899 (0.042) 1.841 (0.044) 1.899 (0.042)
MCC 0.789 (0.051) 0.806 (0.050) 0.197 (0.062) 0.760 (0.050) 0.863 (0.041) 0.901 (0.040) 0.843 (0.043) 0.901 (0.040)
Table 10: Performance results in terms of mean (standard deviation) of the compared methods in scenario S9 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.037 (0.023) 0.033 (0.020) 0.036 (0.021) 0.038 (0.024) 0.060 (0.027) 0.030 (0.020) 0.038 (0.023) 0.031 (0.021)
FPR 0.025 (0.019) 0.022 (0.019) 0.024 (0.020) 0.028 (0.020) 0.037 (0.019) 0.021 (0.020) 0.026 (0.019) 0.022 (0.020)
FNR 0.188 (0.239) 0.147 (0.202) 0.192 (0.243) 0.165 (0.228) 0.478 (0.340) 0.125 (0.192) 0.197 (0.240) 0.130 (0.199)
Accuracy 0.963 (0.023) 0.967 (0.020) 0.964 (0.021) 0.962 (0.024) 0.940 (0.027) 0.970 (0.020) 0.962 (0.023) 0.969 (0.021)
F1 1.600 (0.275) 1.641 (0.271) 1.607 (0.283) 1.602 (0.276) 1.382 (0.296) 1.660 (0.276) 1.593 (0.271) 1.656 (0.279)
MCC 0.685 (0.198) 0.730 (0.173) 0.691 (0.209) 0.689 (0.195) 0.437 (0.289) 0.751 (0.171) 0.677 (0.196) 0.747 (0.178)
$n=100$
ERR 0.029 (0.016) 0.024 (0.015) 0.029 (0.015) 0.029 (0.016) 0.042 (0.018) 0.020 (0.014) 0.028 (0.016) 0.020 (0.014)
FPR 0.015 (0.012) 0.013 (0.012) 0.015 (0.013) 0.018 (0.013) 0.022 (0.013) 0.011 (0.012) 0.015 (0.012) 0.011 (0.012)
FNR 0.230 (0.210) 0.186 (0.188) 0.242 (0.192) 0.186 (0.188) 0.393 (0.224) 0.141 (0.172) 0.227 (0.203) 0.143 (0.175)
Accuracy 0.971 (0.016) 0.976 (0.015) 0.971 (0.015) 0.971 (0.016) 0.958 (0.018) 0.980 (0.014) 0.972 (0.016) 0.980 (0.014)
F1 1.684 (0.186) 1.733 (0.180) 1.683 (0.188) 1.697 (0.182) 1.538 (0.208) 1.775 (0.172) 1.689 (0.180) 1.773 (0.172)
MCC 0.707 (0.169) 0.756 (0.158) 0.705 (0.172) 0.720 (0.161) 0.557 (0.203) 0.799 (0.142) 0.712 (0.162) 0.797 (0.143)
$n=500$
ERR 0.017 (0.006) 0.016 (0.005) 0.024 (0.007) 0.018 (0.006) 0.019 (0.007) 0.009 (0.006) 0.015 (0.006) 0.009 (0.006)
FPR 0.009 (0.005) 0.008 (0.006) 0.012 (0.007) 0.011 (0.006) 0.010 (0.006) 0.005 (0.006) 0.008 (0.005) 0.005 (0.006)
FNR 0.156 (0.092) 0.142 (0.084) 0.226 (0.080) 0.142 (0.090) 0.178 (0.085) 0.069 (0.088) 0.137 (0.089) 0.071 (0.088)
Accuracy 0.983 (0.006) 0.984 (0.005) 0.976 (0.007) 0.982 (0.006) 0.981 (0.007) 0.991 (0.006) 0.985 (0.006) 0.991 (0.006)
F1 1.817 (0.063) 1.832 (0.061) 1.748 (0.082) 1.810 (0.066) 1.797 (0.079) 1.907 (0.061) 1.837 (0.061) 1.906 (0.060)
MCC 0.822 (0.060) 0.837 (0.058) 0.752 (0.081) 0.815 (0.063) 0.801 (0.077) 0.912 (0.055) 0.841 (0.058) 0.910 (0.055)
$n=1000$
ERR 0.015 (0.005) 0.013 (0.004) 0.023 (0.005) 0.015 (0.004) 0.012 (0.004) 0.006 (0.004) 0.012 (0.004) 0.006 (0.004)
FPR 0.008 (0.004) 0.007 (0.004) 0.012 (0.005) 0.009 (0.004) 0.007 (0.004) 0.003 (0.004) 0.006 (0.004) 0.003 (0.004)
FNR 0.144 (0.068) 0.128 (0.061) 0.229 (0.057) 0.127 (0.068) 0.117 (0.065) 0.051 (0.065) 0.111 (0.064) 0.052 (0.065)
Accuracy 0.985 (0.005) 0.987 (0.004) 0.977 (0.005) 0.985 (0.004) 0.988 (0.004) 0.994 (0.004) 0.988 (0.004) 0.994 (0.004)
F1 1.840 (0.049) 1.857 (0.042) 1.754 (0.062) 1.842 (0.045) 1.868 (0.048) 1.937 (0.044) 1.874 (0.041) 1.937 (0.044)
MCC 0.842 (0.048) 0.860 (0.040) 0.756 (0.061) 0.844 (0.043) 0.871 (0.046) 0.940 (0.040) 0.877 (0.039) 0.939 (0.040)
Table 11: Performance results in terms of mean (standard deviation) of the compared methods in scenario S10 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.044 (0.026) 0.039 (0.022) 0.058 (0.027) 0.044 (0.025) 0.041 (0.023) 0.037 (0.023) 0.041 (0.023) 0.036 (0.022)
FPR 0.029 (0.019) 0.026 (0.019) 0.037 (0.019) 0.031 (0.022) 0.027 (0.019) 0.025 (0.019) 0.027 (0.019) 0.024 (0.019)
FNR 0.252 (0.288) 0.213 (0.260) 0.459 (0.341) 0.248 (0.280) 0.229 (0.266) 0.194 (0.251) 0.238 (0.272) 0.174 (0.235)
Accuracy 0.956 (0.026) 0.961 (0.022) 0.942 (0.027) 0.956 (0.025) 0.959 (0.023) 0.963 (0.023) 0.959 (0.023) 0.964 (0.022)
F1 1.537 (0.290) 1.577 (0.286) 1.389 (0.298) 1.537 (0.294) 1.560 (0.283) 1.591 (0.288) 1.557 (0.287) 1.608 (0.286)
MCC 0.622 (0.233) 0.667 (0.213) 0.452 (0.287) 0.623 (0.235) 0.647 (0.216) 0.683 (0.209) 0.644 (0.223) 0.702 (0.197)
$n=100$
ERR 0.031 (0.016) 0.028 (0.015) 0.051 (0.019) 0.031 (0.016) 0.029 (0.015) 0.024 (0.015) 0.029 (0.015) 0.024 (0.015)
FPR 0.017 (0.012) 0.015 (0.012) 0.027 (0.013) 0.018 (0.014) 0.016 (0.012) 0.013 (0.012) 0.016 (0.013) 0.013 (0.012)
FNR 0.249 (0.206) 0.223 (0.201) 0.487 (0.244) 0.222 (0.199) 0.227 (0.199) 0.181 (0.188) 0.231 (0.200) 0.174 (0.184)
Accuracy 0.969 (0.016) 0.972 (0.015) 0.949 (0.019) 0.969 (0.016) 0.971 (0.015) 0.976 (0.015) 0.971 (0.015) 0.976 (0.015)
F1 1.659 (0.185) 1.687 (0.186) 1.443 (0.219) 1.667 (0.194) 1.681 (0.181) 1.727 (0.180) 1.679 (0.188) 1.732 (0.179)
MCC 0.682 (0.168) 0.710 (0.168) 0.459 (0.218) 0.692 (0.175) 0.704 (0.162) 0.751 (0.156) 0.702 (0.170) 0.757 (0.154)
$n=500$
ERR 0.018 (0.007) 0.016 (0.006) 0.053 (0.009) 0.017 (0.006) 0.015 (0.006) 0.010 (0.006) 0.016 (0.006) 0.010 (0.006)
FPR 0.009 (0.006) 0.008 (0.006) 0.028 (0.006) 0.009 (0.006) 0.008 (0.005) 0.005 (0.006) 0.008 (0.005) 0.005 (0.006)
FNR 0.161 (0.094) 0.144 (0.092) 0.530 (0.093) 0.145 (0.090) 0.132 (0.091) 0.088 (0.089) 0.141 (0.091) 0.088 (0.089)
Accuracy 0.982 (0.007) 0.984 (0.006) 0.947 (0.009) 0.983 (0.006) 0.985 (0.006) 0.990 (0.006) 0.984 (0.006) 0.990 (0.006)
F1 1.814 (0.070) 1.832 (0.067) 1.436 (0.099) 1.822 (0.067) 1.845 (0.065) 1.890 (0.064) 1.835 (0.065) 1.890 (0.064)
MCC 0.818 (0.067) 0.837 (0.064) 0.439 (0.099) 0.827 (0.063) 0.849 (0.061) 0.895 (0.058) 0.840 (0.062) 0.895 (0.058)
$n=1000$
ERR 0.014 (0.004) 0.013 (0.004) 0.054 (0.007) 0.013 (0.004) 0.011 (0.004) 0.007 (0.004) 0.012 (0.004) 0.007 (0.004)
FPR 0.007 (0.004) 0.007 (0.004) 0.029 (0.004) 0.007 (0.004) 0.006 (0.004) 0.004 (0.004) 0.006 (0.004) 0.004 (0.004)
FNR 0.133 (0.066) 0.118 (0.068) 0.541 (0.069) 0.120 (0.068) 0.103 (0.066) 0.065 (0.065) 0.113 (0.067) 0.065 (0.065)
Accuracy 0.986 (0.004) 0.987 (0.004) 0.946 (0.007) 0.987 (0.004) 0.989 (0.004) 0.993 (0.004) 0.988 (0.004) 0.993 (0.004)
F1 1.852 (0.045) 1.868 (0.044) 1.427 (0.071) 1.859 (0.044) 1.883 (0.042) 1.923 (0.041) 1.873 (0.043) 1.923 (0.041)
MCC 0.854 (0.044) 0.870 (0.042) 0.429 (0.071) 0.861 (0.043) 0.885 (0.040) 0.925 (0.038) 0.875 (0.041) 0.925 (0.038)
Table 12: Performance results in terms of mean (standard deviation) of the compared methods in scenario S11 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.047 (0.027) 0.041 (0.024) 0.072 (0.029) 0.044 (0.026) 0.043 (0.024) 0.040 (0.024) 0.040 (0.023) 0.038 (0.024)
FPR 0.030 (0.019) 0.027 (0.019) 0.043 (0.017) 0.031 (0.024) 0.028 (0.019) 0.027 (0.020) 0.027 (0.019) 0.026 (0.020)
FNR 0.287 (0.306) 0.234 (0.277) 0.609 (0.339) 0.231 (0.268) 0.253 (0.280) 0.226 (0.269) 0.226 (0.262) 0.203 (0.255)
Accuracy 0.953 (0.027) 0.959 (0.024) 0.928 (0.029) 0.956 (0.026) 0.957 (0.024) 0.960 (0.024) 0.960 (0.023) 0.962 (0.024)
F1 1.512 (0.296) 1.561 (0.294) 1.271 (0.284) 1.548 (0.293) 1.544 (0.289) 1.566 (0.293) 1.567 (0.287) 1.585 (0.290)
MCC 0.595 (0.249) 0.650 (0.228) 0.317 (0.289) 0.639 (0.226) 0.632 (0.227) 0.657 (0.223) 0.657 (0.215) 0.678 (0.212)
$n=100$
ERR 0.033 (0.017) 0.028 (0.015) 0.059 (0.020) 0.030 (0.017) 0.029 (0.015) 0.026 (0.015) 0.027 (0.015) 0.026 (0.015)
FPR 0.018 (0.012) 0.015 (0.012) 0.031 (0.012) 0.018 (0.014) 0.015 (0.012) 0.014 (0.012) 0.015 (0.012) 0.014 (0.012)
FNR 0.280 (0.217) 0.228 (0.194) 0.582 (0.238) 0.213 (0.195) 0.231 (0.193) 0.205 (0.188) 0.215 (0.193) 0.196 (0.185)
Accuracy 0.967 (0.017) 0.972 (0.015) 0.941 (0.020) 0.970 (0.017) 0.971 (0.015) 0.974 (0.015) 0.973 (0.015) 0.974 (0.015)
F1 1.636 (0.197) 1.687 (0.183) 1.358 (0.214) 1.679 (0.194) 1.683 (0.180) 1.709 (0.181) 1.698 (0.180) 1.717 (0.177)
MCC 0.658 (0.182) 0.711 (0.162) 0.372 (0.216) 0.704 (0.172) 0.707 (0.159) 0.734 (0.157) 0.722 (0.158) 0.742 (0.152)
$n=500$
ERR 0.019 (0.007) 0.016 (0.006) 0.061 (0.009) 0.018 (0.006) 0.015 (0.006) 0.011 (0.006) 0.015 (0.006) 0.011 (0.006)
FPR 0.010 (0.005) 0.008 (0.005) 0.032 (0.006) 0.011 (0.006) 0.008 (0.005) 0.006 (0.005) 0.008 (0.005) 0.006 (0.005)
FNR 0.175 (0.095) 0.144 (0.093) 0.608 (0.092) 0.150 (0.093) 0.139 (0.093) 0.096 (0.089) 0.136 (0.093) 0.096 (0.089)
Accuracy 0.981 (0.007) 0.984 (0.006) 0.939 (0.009) 0.982 (0.006) 0.985 (0.006) 0.989 (0.006) 0.985 (0.006) 0.989 (0.006)
F1 1.800 (0.068) 1.832 (0.064) 1.355 (0.097) 1.810 (0.065) 1.837 (0.061) 1.882 (0.059) 1.841 (0.061) 1.883 (0.059)
MCC 0.804 (0.066) 0.836 (0.061) 0.357 (0.097) 0.815 (0.063) 0.842 (0.058) 0.887 (0.055) 0.845 (0.058) 0.887 (0.054)
$n=1000$
ERR 0.015 (0.004) 0.013 (0.004) 0.061 (0.007) 0.015 (0.004) 0.012 (0.004) 0.008 (0.004) 0.012 (0.004) 0.008 (0.004)
FPR 0.008 (0.004) 0.007 (0.004) 0.032 (0.004) 0.009 (0.004) 0.006 (0.004) 0.004 (0.004) 0.006 (0.004) 0.004 (0.004)
FNR 0.147 (0.064) 0.119 (0.068) 0.608 (0.065) 0.132 (0.069) 0.110 (0.066) 0.071 (0.065) 0.107 (0.067) 0.070 (0.065)
Accuracy 0.985 (0.004) 0.987 (0.004) 0.939 (0.007) 0.985 (0.004) 0.988 (0.004) 0.992 (0.004) 0.988 (0.004) 0.992 (0.004)
F1 1.836 (0.046) 1.865 (0.042) 1.357 (0.067) 1.839 (0.045) 1.874 (0.040) 1.916 (0.041) 1.878 (0.040) 1.916 (0.041)
MCC 0.839 (0.045) 0.867 (0.041) 0.358 (0.067) 0.841 (0.044) 0.877 (0.038) 0.919 (0.038) 0.880 (0.038) 0.919 (0.038)
Table 13: Performance results in terms of mean (standard deviation) of the compared methods in scenario S12 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.055 (0.028) 0.052 (0.027) 0.084 (0.031) 0.059 (0.029) 0.045 (0.024) 0.046 (0.026) 0.047 (0.025) 0.045 (0.025)
FPR 0.035 (0.018) 0.033 (0.018) 0.050 (0.015) 0.041 (0.024) 0.029 (0.019) 0.030 (0.019) 0.030 (0.019) 0.029 (0.019)
FNR 0.385 (0.332) 0.357 (0.322) 0.732 (0.321) 0.351 (0.318) 0.275 (0.287) 0.293 (0.306) 0.302 (0.301) 0.271 (0.294)
Accuracy 0.945 (0.028) 0.948 (0.027) 0.916 (0.031) 0.941 (0.029) 0.955 (0.024) 0.954 (0.026) 0.953 (0.025) 0.955 (0.025)
F1 1.429 (0.286) 1.455 (0.289) 1.162 (0.251) 1.430 (0.280) 1.521 (0.283) 1.508 (0.294) 1.503 (0.291) 1.525 (0.291)
MCC 0.500 (0.262) 0.530 (0.258) 0.194 (0.266) 0.507 (0.249) 0.605 (0.230) 0.590 (0.249) 0.585 (0.246) 0.610 (0.239)
$n=100$
ERR 0.041 (0.019) 0.036 (0.018) 0.073 (0.022) 0.043 (0.020) 0.031 (0.016) 0.029 (0.015) 0.032 (0.016) 0.029 (0.015)
FPR 0.021 (0.012) 0.019 (0.013) 0.038 (0.011) 0.026 (0.015) 0.016 (0.012) 0.015 (0.012) 0.016 (0.012) 0.015 (0.012)
FNR 0.372 (0.231) 0.323 (0.220) 0.715 (0.217) 0.328 (0.228) 0.265 (0.203) 0.245 (0.198) 0.270 (0.204) 0.240 (0.195)
Accuracy 0.959 (0.019) 0.964 (0.018) 0.927 (0.022) 0.957 (0.020) 0.969 (0.016) 0.971 (0.015) 0.968 (0.016) 0.971 (0.015)
F1 1.559 (0.207) 1.608 (0.198) 1.229 (0.194) 1.561 (0.202) 1.663 (0.181) 1.683 (0.180) 1.658 (0.187) 1.688 (0.180)
MCC 0.577 (0.202) 0.627 (0.191) 0.238 (0.199) 0.581 (0.195) 0.683 (0.170) 0.703 (0.166) 0.678 (0.175) 0.708 (0.165)
$n=500$
ERR 0.024 (0.007) 0.019 (0.006) 0.070 (0.009) 0.026 (0.007) 0.017 (0.006) 0.012 (0.006) 0.016 (0.006) 0.012 (0.006)
FPR 0.013 (0.006) 0.010 (0.005) 0.037 (0.005) 0.016 (0.006) 0.009 (0.006) 0.006 (0.006) 0.008 (0.005) 0.006 (0.006)
FNR 0.226 (0.092) 0.174 (0.092) 0.701 (0.086) 0.212 (0.098) 0.157 (0.087) 0.108 (0.087) 0.146 (0.091) 0.108 (0.086)
Accuracy 0.976 (0.007) 0.981 (0.006) 0.930 (0.009) 0.974 (0.007) 0.983 (0.006) 0.988 (0.006) 0.984 (0.006) 0.988 (0.006)
F1 1.748 (0.079) 1.801 (0.068) 1.259 (0.090) 1.734 (0.075) 1.819 (0.065) 1.869 (0.063) 1.830 (0.066) 1.870 (0.063)
MCC 0.752 (0.077) 0.805 (0.066) 0.261 (0.090) 0.739 (0.073) 0.823 (0.062) 0.874 (0.059) 0.834 (0.063) 0.874 (0.058)
$n=1000$
ERR 0.020 (0.005) 0.015 (0.004) 0.070 (0.007) 0.022 (0.005) 0.013 (0.004) 0.009 (0.004) 0.013 (0.004) 0.009 (0.004)
FPR 0.010 (0.004) 0.008 (0.004) 0.037 (0.004) 0.013 (0.004) 0.007 (0.004) 0.005 (0.004) 0.007 (0.004) 0.004 (0.004)
FNR 0.189 (0.064) 0.139 (0.068) 0.696 (0.060) 0.190 (0.072) 0.127 (0.064) 0.078 (0.065) 0.118 (0.065) 0.077 (0.065)
Accuracy 0.980 (0.005) 0.985 (0.004) 0.930 (0.007) 0.978 (0.005) 0.987 (0.004) 0.991 (0.004) 0.987 (0.004) 0.991 (0.004)
F1 1.793 (0.051) 1.845 (0.045) 1.265 (0.063) 1.772 (0.052) 1.858 (0.043) 1.909 (0.041) 1.867 (0.042) 1.910 (0.041)
MCC 0.795 (0.050) 0.848 (0.044) 0.266 (0.064) 0.774 (0.051) 0.860 (0.042) 0.912 (0.039) 0.869 (0.041) 0.912 (0.039)
Table 14: Performance results in terms of mean (standard deviation) of the compared methods in scenario S13 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.037 (0.021) 0.033 (0.021) 0.036 (0.021) 0.036 (0.022) 0.055 (0.026) 0.033 (0.021) 0.037 (0.022) 0.033 (0.021)
FPR 0.025 (0.019) 0.023 (0.020) 0.024 (0.020) 0.026 (0.021) 0.034 (0.020) 0.022 (0.020) 0.025 (0.019) 0.022 (0.020)
FNR 0.185 (0.233) 0.161 (0.221) 0.198 (0.246) 0.167 (0.235) 0.433 (0.337) 0.150 (0.218) 0.192 (0.242) 0.151 (0.216)
Accuracy 0.963 (0.021) 0.967 (0.021) 0.964 (0.021) 0.964 (0.022) 0.945 (0.026) 0.967 (0.021) 0.963 (0.022) 0.967 (0.021)
F1 1.606 (0.270) 1.632 (0.278) 1.604 (0.284) 1.613 (0.283) 1.421 (0.301) 1.640 (0.280) 1.600 (0.275) 1.639 (0.279)
MCC 0.690 (0.190) 0.718 (0.192) 0.687 (0.213) 0.698 (0.206) 0.480 (0.287) 0.727 (0.192) 0.683 (0.201) 0.726 (0.189)
$n=100$
ERR 0.028 (0.016) 0.025 (0.014) 0.031 (0.015) 0.028 (0.015) 0.041 (0.018) 0.021 (0.014) 0.028 (0.015) 0.022 (0.014)
FPR 0.015 (0.012) 0.013 (0.013) 0.016 (0.013) 0.017 (0.013) 0.022 (0.013) 0.011 (0.012) 0.015 (0.012) 0.012 (0.013)
FNR 0.213 (0.200) 0.192 (0.183) 0.265 (0.202) 0.187 (0.184) 0.379 (0.227) 0.148 (0.170) 0.220 (0.196) 0.155 (0.172)
Accuracy 0.972 (0.016) 0.975 (0.014) 0.969 (0.015) 0.973 (0.015) 0.959 (0.018) 0.979 (0.014) 0.972 (0.015) 0.978 (0.014)
F1 1.693 (0.179) 1.721 (0.177) 1.657 (0.197) 1.701 (0.185) 1.550 (0.213) 1.761 (0.176) 1.688 (0.178) 1.755 (0.179)
MCC 0.718 (0.158) 0.746 (0.153) 0.680 (0.183) 0.725 (0.162) 0.569 (0.207) 0.787 (0.145) 0.713 (0.158) 0.781 (0.150)
$n=500$
ERR 0.017 (0.006) 0.016 (0.006) 0.032 (0.008) 0.017 (0.006) 0.019 (0.006) 0.009 (0.006) 0.016 (0.006) 0.010 (0.006)
FPR 0.009 (0.005) 0.008 (0.006) 0.017 (0.006) 0.010 (0.006) 0.010 (0.006) 0.005 (0.006) 0.008 (0.006) 0.005 (0.006)
FNR 0.153 (0.093) 0.148 (0.083) 0.314 (0.089) 0.141 (0.085) 0.178 (0.084) 0.080 (0.088) 0.149 (0.086) 0.081 (0.087)
Accuracy 0.983 (0.006) 0.984 (0.006) 0.968 (0.008) 0.983 (0.006) 0.981 (0.006) 0.991 (0.006) 0.984 (0.006) 0.990 (0.006)
F1 1.823 (0.065) 1.830 (0.062) 1.658 (0.083) 1.822 (0.065) 1.799 (0.071) 1.900 (0.065) 1.828 (0.063) 1.899 (0.064)
MCC 0.828 (0.062) 0.835 (0.059) 0.662 (0.082) 0.827 (0.061) 0.804 (0.069) 0.905 (0.059) 0.833 (0.059) 0.904 (0.059)
$n=1000$
ERR 0.015 (0.004) 0.013 (0.004) 0.034 (0.006) 0.014 (0.004) 0.013 (0.004) 0.007 (0.004) 0.013 (0.004) 0.007 (0.004)
FPR 0.008 (0.004) 0.007 (0.004) 0.018 (0.004) 0.008 (0.004) 0.007 (0.004) 0.003 (0.004) 0.007 (0.004) 0.004 (0.004)
FNR 0.139 (0.067) 0.127 (0.060) 0.337 (0.069) 0.123 (0.063) 0.125 (0.064) 0.061 (0.066) 0.122 (0.064) 0.062 (0.066)
Accuracy 0.985 (0.004) 0.987 (0.004) 0.966 (0.006) 0.986 (0.004) 0.987 (0.004) 0.993 (0.004) 0.987 (0.004) 0.993 (0.004)
F1 1.847 (0.046) 1.860 (0.040) 1.641 (0.064) 1.853 (0.041) 1.861 (0.041) 1.928 (0.041) 1.865 (0.040) 1.927 (0.041)
MCC 0.849 (0.045) 0.862 (0.039) 0.642 (0.063) 0.855 (0.040) 0.863 (0.040) 0.930 (0.039) 0.867 (0.039) 0.929 (0.039)
Table 15: Performance results in terms of mean (standard deviation) of the compared methods in scenario S14 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.046 (0.025) 0.047 (0.025) 0.065 (0.028) 0.051 (0.026) 0.045 (0.024) 0.038 (0.023) 0.048 (0.025) 0.039 (0.024)
FPR 0.029 (0.019) 0.030 (0.019) 0.039 (0.017) 0.034 (0.021) 0.029 (0.019) 0.025 (0.019) 0.031 (0.019) 0.025 (0.020)
FNR 0.301 (0.298) 0.304 (0.297) 0.517 (0.338) 0.342 (0.310) 0.288 (0.292) 0.208 (0.258) 0.323 (0.307) 0.223 (0.271)
Accuracy 0.954 (0.025) 0.953 (0.025) 0.935 (0.028) 0.949 (0.026) 0.955 (0.024) 0.962 (0.023) 0.952 (0.025) 0.961 (0.024)
F1 1.518 (0.284) 1.514 (0.281) 1.344 (0.278) 1.480 (0.288) 1.529 (0.280) 1.596 (0.283) 1.501 (0.287) 1.586 (0.292)
MCC 0.588 (0.245) 0.585 (0.242) 0.393 (0.276) 0.547 (0.258) 0.601 (0.237) 0.675 (0.217) 0.570 (0.253) 0.665 (0.233)
$n=100$
ERR 0.033 (0.017) 0.035 (0.017) 0.061 (0.021) 0.040 (0.018) 0.032 (0.016) 0.025 (0.015) 0.038 (0.017) 0.025 (0.015)
FPR 0.017 (0.012) 0.018 (0.012) 0.032 (0.012) 0.023 (0.014) 0.017 (0.012) 0.013 (0.012) 0.020 (0.012) 0.013 (0.012)
FNR 0.276 (0.211) 0.300 (0.212) 0.586 (0.225) 0.325 (0.210) 0.265 (0.205) 0.182 (0.186) 0.332 (0.209) 0.193 (0.188)
Accuracy 0.967 (0.017) 0.965 (0.017) 0.939 (0.021) 0.960 (0.018) 0.968 (0.016) 0.975 (0.015) 0.962 (0.017) 0.975 (0.015)
F1 1.642 (0.177) 1.620 (0.182) 1.350 (0.203) 1.580 (0.194) 1.654 (0.177) 1.733 (0.163) 1.592 (0.188) 1.724 (0.165)
MCC 0.662 (0.168) 0.640 (0.174) 0.362 (0.205) 0.599 (0.187) 0.674 (0.167) 0.755 (0.144) 0.611 (0.181) 0.746 (0.148)
$n=500$
ERR 0.020 (0.007) 0.022 (0.007) 0.064 (0.010) 0.025 (0.007) 0.017 (0.006) 0.011 (0.006) 0.023 (0.007) 0.011 (0.006)
FPR 0.011 (0.006) 0.011 (0.006) 0.034 (0.005) 0.013 (0.006) 0.009 (0.006) 0.006 (0.006) 0.012 (0.006) 0.006 (0.006)
FNR 0.189 (0.094) 0.204 (0.092) 0.639 (0.093) 0.227 (0.088) 0.159 (0.091) 0.094 (0.089) 0.224 (0.086) 0.096 (0.088)
Accuracy 0.980 (0.007) 0.978 (0.007) 0.936 (0.010) 0.975 (0.007) 0.983 (0.006) 0.989 (0.006) 0.977 (0.007) 0.989 (0.006)
F1 1.785 (0.073) 1.770 (0.073) 1.320 (0.094) 1.741 (0.079) 1.817 (0.067) 1.884 (0.063) 1.751 (0.076) 1.881 (0.062)
MCC 0.790 (0.070) 0.774 (0.071) 0.323 (0.094) 0.745 (0.077) 0.821 (0.064) 0.888 (0.058) 0.755 (0.074) 0.886 (0.057)
$n=1000$
ERR 0.017 (0.005) 0.017 (0.004) 0.063 (0.007) 0.020 (0.005) 0.013 (0.004) 0.008 (0.004) 0.019 (0.004) 0.008 (0.004)
FPR 0.009 (0.004) 0.009 (0.004) 0.033 (0.004) 0.011 (0.004) 0.007 (0.004) 0.004 (0.004) 0.010 (0.004) 0.004 (0.004)
FNR 0.162 (0.067) 0.167 (0.066) 0.631 (0.068) 0.183 (0.061) 0.126 (0.065) 0.069 (0.065) 0.183 (0.060) 0.071 (0.065)
Accuracy 0.983 (0.005) 0.983 (0.004) 0.937 (0.007) 0.980 (0.005) 0.987 (0.004) 0.992 (0.004) 0.981 (0.004) 0.992 (0.004)
F1 1.821 (0.050) 1.816 (0.047) 1.333 (0.069) 1.794 (0.051) 1.859 (0.045) 1.918 (0.041) 1.800 (0.050) 1.917 (0.041)
MCC 0.823 (0.049) 0.818 (0.046) 0.334 (0.069) 0.796 (0.050) 0.861 (0.044) 0.920 (0.038) 0.802 (0.049) 0.919 (0.039)
Table 16: Performance results in terms of mean (standard deviation) of the compared methods in scenario S15 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.049 (0.027) 0.048 (0.026) 0.075 (0.031) 0.054 (0.027) 0.046 (0.025) 0.042 (0.024) 0.049 (0.025) 0.042 (0.024)
FPR 0.031 (0.019) 0.031 (0.019) 0.045 (0.017) 0.036 (0.023) 0.030 (0.018) 0.027 (0.019) 0.031 (0.019) 0.027 (0.019)
FNR 0.333 (0.315) 0.324 (0.308) 0.629 (0.343) 0.337 (0.315) 0.290 (0.294) 0.249 (0.277) 0.335 (0.307) 0.250 (0.278)
Accuracy 0.951 (0.027) 0.952 (0.026) 0.925 (0.031) 0.947 (0.027) 0.954 (0.025) 0.958 (0.024) 0.951 (0.025) 0.958 (0.024)
F1 1.487 (0.292) 1.496 (0.288) 1.249 (0.276) 1.468 (0.291) 1.520 (0.282) 1.556 (0.286) 1.488 (0.287) 1.557 (0.289)
MCC 0.559 (0.259) 0.569 (0.251) 0.289 (0.285) 0.541 (0.259) 0.596 (0.236) 0.637 (0.228) 0.560 (0.252) 0.638 (0.232)
$n=100$
ERR 0.037 (0.018) 0.037 (0.018) 0.067 (0.022) 0.040 (0.018) 0.033 (0.017) 0.028 (0.016) 0.037 (0.017) 0.028 (0.016)
FPR 0.019 (0.012) 0.019 (0.012) 0.035 (0.012) 0.023 (0.014) 0.017 (0.012) 0.014 (0.012) 0.019 (0.013) 0.014 (0.012)
FNR 0.319 (0.213) 0.328 (0.219) 0.648 (0.236) 0.321 (0.218) 0.274 (0.211) 0.224 (0.200) 0.322 (0.216) 0.229 (0.204)
Accuracy 0.963 (0.018) 0.963 (0.018) 0.933 (0.022) 0.960 (0.018) 0.967 (0.017) 0.972 (0.016) 0.963 (0.017) 0.972 (0.016)
F1 1.603 (0.188) 1.597 (0.196) 1.290 (0.206) 1.581 (0.198) 1.647 (0.188) 1.697 (0.182) 1.603 (0.195) 1.693 (0.184)
MCC 0.624 (0.178) 0.618 (0.186) 0.302 (0.210) 0.602 (0.188) 0.669 (0.174) 0.720 (0.164) 0.624 (0.185) 0.716 (0.167)
$n=500$
ERR 0.023 (0.007) 0.023 (0.007) 0.068 (0.010) 0.024 (0.007) 0.018 (0.006) 0.012 (0.006) 0.022 (0.007) 0.012 (0.006)
FPR 0.012 (0.006) 0.012 (0.006) 0.036 (0.005) 0.014 (0.006) 0.009 (0.006) 0.006 (0.006) 0.012 (0.006) 0.006 (0.006)
FNR 0.221 (0.094) 0.217 (0.094) 0.681 (0.090) 0.220 (0.092) 0.162 (0.091) 0.105 (0.088) 0.211 (0.090) 0.107 (0.088)
Accuracy 0.977 (0.007) 0.977 (0.007) 0.932 (0.010) 0.976 (0.007) 0.982 (0.006) 0.988 (0.006) 0.978 (0.007) 0.988 (0.006)
F1 1.753 (0.078) 1.756 (0.076) 1.278 (0.089) 1.743 (0.078) 1.813 (0.067) 1.872 (0.061) 1.763 (0.075) 1.871 (0.062)
MCC 0.757 (0.076) 0.760 (0.074) 0.279 (0.089) 0.748 (0.076) 0.818 (0.064) 0.877 (0.056) 0.768 (0.073) 0.875 (0.058)
$n=1000$
ERR 0.020 (0.005) 0.019 (0.004) 0.067 (0.007) 0.020 (0.005) 0.014 (0.004) 0.008 (0.004) 0.018 (0.004) 0.008 (0.004)
FPR 0.010 (0.004) 0.010 (0.004) 0.035 (0.004) 0.011 (0.004) 0.007 (0.004) 0.004 (0.004) 0.009 (0.004) 0.004 (0.004)
FNR 0.192 (0.066) 0.179 (0.065) 0.666 (0.066) 0.181 (0.065) 0.130 (0.066) 0.077 (0.066) 0.171 (0.061) 0.078 (0.065)
Accuracy 0.980 (0.005) 0.981 (0.004) 0.933 (0.007) 0.980 (0.005) 0.986 (0.004) 0.992 (0.004) 0.982 (0.004) 0.992 (0.004)
F1 1.791 (0.051) 1.804 (0.047) 1.297 (0.067) 1.793 (0.049) 1.856 (0.044) 1.911 (0.042) 1.813 (0.049) 1.911 (0.042)
MCC 0.793 (0.051) 0.807 (0.046) 0.298 (0.067) 0.795 (0.048) 0.859 (0.042) 0.914 (0.040) 0.816 (0.048) 0.913 (0.039)
Table 17: Performance results in terms of mean (standard deviation) of the compared methods in scenario S16 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.062 (0.030) 0.058 (0.029) 0.086 (0.032) 0.069 (0.032) 0.050 (0.026) 0.048 (0.027) 0.055 (0.028) 0.048 (0.027)
FPR 0.038 (0.018) 0.036 (0.018) 0.051 (0.015) 0.046 (0.025) 0.032 (0.018) 0.031 (0.019) 0.035 (0.019) 0.030 (0.019)
FNR 0.471 (0.343) 0.428 (0.341) 0.754 (0.310) 0.452 (0.342) 0.341 (0.305) 0.318 (0.310) 0.400 (0.331) 0.316 (0.311)
Accuracy 0.938 (0.030) 0.942 (0.029) 0.914 (0.032) 0.931 (0.032) 0.950 (0.026) 0.952 (0.027) 0.945 (0.028) 0.952 (0.027)
F1 1.371 (0.288) 1.406 (0.291) 1.148 (0.246) 1.359 (0.284) 1.478 (0.281) 1.497 (0.294) 1.432 (0.293) 1.500 (0.297)
MCC 0.430 (0.279) 0.469 (0.276) 0.176 (0.261) 0.422 (0.274) 0.551 (0.245) 0.573 (0.256) 0.499 (0.271) 0.576 (0.259)
$n=100$
ERR 0.048 (0.020) 0.044 (0.019) 0.076 (0.022) 0.052 (0.021) 0.036 (0.017) 0.032 (0.017) 0.040 (0.018) 0.032 (0.017)
FPR 0.025 (0.012) 0.023 (0.013) 0.039 (0.010) 0.031 (0.016) 0.018 (0.012) 0.016 (0.013) 0.021 (0.012) 0.017 (0.013)
FNR 0.439 (0.235) 0.409 (0.235) 0.749 (0.213) 0.407 (0.235) 0.313 (0.211) 0.274 (0.211) 0.361 (0.221) 0.283 (0.212)
Accuracy 0.952 (0.020) 0.956 (0.019) 0.924 (0.022) 0.948 (0.021) 0.964 (0.017) 0.968 (0.017) 0.960 (0.018) 0.968 (0.017)
F1 1.491 (0.205) 1.523 (0.210) 1.198 (0.192) 1.485 (0.206) 1.615 (0.195) 1.654 (0.196) 1.569 (0.199) 1.648 (0.199)
MCC 0.509 (0.201) 0.542 (0.205) 0.206 (0.196) 0.506 (0.199) 0.637 (0.182) 0.676 (0.180) 0.589 (0.190) 0.669 (0.184)
$n=500$
ERR 0.029 (0.007) 0.027 (0.008) 0.072 (0.010) 0.032 (0.008) 0.021 (0.007) 0.014 (0.006) 0.024 (0.007) 0.014 (0.006)
FPR 0.015 (0.006) 0.014 (0.006) 0.038 (0.005) 0.019 (0.007) 0.011 (0.006) 0.007 (0.006) 0.013 (0.006) 0.008 (0.006)
FNR 0.283 (0.092) 0.256 (0.093) 0.720 (0.088) 0.276 (0.095) 0.197 (0.089) 0.126 (0.089) 0.226 (0.088) 0.129 (0.089)
Accuracy 0.971 (0.007) 0.973 (0.008) 0.928 (0.010) 0.968 (0.008) 0.979 (0.007) 0.986 (0.006) 0.976 (0.007) 0.986 (0.006)
F1 1.689 (0.080) 1.716 (0.081) 1.238 (0.089) 1.670 (0.083) 1.777 (0.074) 1.850 (0.067) 1.748 (0.077) 1.848 (0.069)
MCC 0.693 (0.079) 0.720 (0.080) 0.239 (0.090) 0.675 (0.082) 0.782 (0.072) 0.855 (0.064) 0.752 (0.075) 0.852 (0.066)
$n=1000$
ERR 0.026 (0.005) 0.022 (0.005) 0.070 (0.007) 0.027 (0.006) 0.017 (0.004) 0.010 (0.004) 0.019 (0.005) 0.010 (0.004)
FPR 0.014 (0.004) 0.012 (0.004) 0.037 (0.003) 0.016 (0.005) 0.009 (0.004) 0.005 (0.004) 0.010 (0.004) 0.005 (0.004)
FNR 0.255 (0.065) 0.215 (0.067) 0.699 (0.062) 0.239 (0.073) 0.162 (0.062) 0.093 (0.067) 0.183 (0.064) 0.094 (0.066)
Accuracy 0.974 (0.005) 0.978 (0.005) 0.930 (0.007) 0.973 (0.006) 0.983 (0.004) 0.990 (0.004) 0.981 (0.005) 0.990 (0.004)
F1 1.725 (0.057) 1.767 (0.053) 1.262 (0.065) 1.722 (0.056) 1.823 (0.050) 1.894 (0.042) 1.800 (0.052) 1.893 (0.043)
MCC 0.727 (0.056) 0.769 (0.053) 0.263 (0.065) 0.724 (0.055) 0.825 (0.049) 0.896 (0.040) 0.802 (0.051) 0.895 (0.041)
Table 18: Performance results in terms of mean (standard deviation) of the compared methods in scenario S17 across the different sample sizes.
Measures w/o copula Measures w copula
Metrics $M_{0}$:KDE $M_{1}$:$k$NN-Eucl $M_{2}$:$k$NN-CDF $M_{3}$:$\epsilon$-CDF $M^{\text{NPCop}}_{0}$:DE $M^{\text{PCop}}_{0}$:DE $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF $M^{\text{PCop}}_{3}$:$\epsilon$-CDF
$n=50$
ERR 0.072 (0.027) 0.053 (0.027) 0.060 (0.026) 0.070 (0.033) 0.049 (0.025) 0.041 (0.024) 0.053 (0.026) 0.041 (0.024)
FPR 0.047 (0.015) 0.037 (0.017) 0.041 (0.017) 0.052 (0.025) 0.035 (0.017) 0.031 (0.019) 0.037 (0.017) 0.030 (0.019)
FNR 0.620 (0.358) 0.343 (0.346) 0.466 (0.360) 0.420 (0.368) 0.295 (0.322) 0.200 (0.276) 0.342 (0.338) 0.200 (0.277)
Accuracy 0.928 (0.027) 0.947 (0.027) 0.940 (0.026) 0.930 (0.033) 0.951 (0.025) 0.959 (0.024) 0.947 (0.026) 0.959 (0.024)
F1 1.229 (0.267) 1.417 (0.291) 1.342 (0.290) 1.321 (0.272) 1.454 (0.289) 1.531 (0.298) 1.422 (0.290) 1.538 (0.300)
MCC 0.279 (0.278) 0.503 (0.263) 0.412 (0.282) 0.399 (0.265) 0.546 (0.247) 0.635 (0.226) 0.508 (0.260) 0.639 (0.230)
$n=100$
ERR 0.064 (0.020) 0.042 (0.020) 0.056 (0.019) 0.052 (0.023) 0.036 (0.018) 0.029 (0.017) 0.038 (0.019) 0.028 (0.017)
FPR 0.037 (0.011) 0.026 (0.011) 0.033 (0.011) 0.035 (0.016) 0.022 (0.012) 0.018 (0.013) 0.023 (0.012) 0.018 (0.013)
FNR 0.646 (0.249) 0.350 (0.258) 0.549 (0.254) 0.372 (0.275) 0.269 (0.234) 0.199 (0.217) 0.307 (0.247) 0.197 (0.219)
Accuracy 0.936 (0.020) 0.958 (0.020) 0.944 (0.019) 0.948 (0.023) 0.964 (0.018) 0.971 (0.017) 0.962 (0.019) 0.972 (0.017)
F1 1.269 (0.206) 1.515 (0.213) 1.354 (0.212) 1.455 (0.218) 1.589 (0.203) 1.661 (0.214) 1.557 (0.212) 1.666 (0.215)
MCC 0.284 (0.210) 0.543 (0.201) 0.373 (0.212) 0.486 (0.209) 0.619 (0.182) 0.693 (0.186) 0.586 (0.196) 0.696 (0.189)
$n=500$
ERR 0.063 (0.009) 0.027 (0.008) 0.057 (0.010) 0.029 (0.008) 0.020 (0.007) 0.019 (0.007) 0.022 (0.007) 0.019 (0.007)
FPR 0.037 (0.005) 0.018 (0.006) 0.034 (0.006) 0.019 (0.006) 0.014 (0.006) 0.014 (0.006) 0.015 (0.006) 0.013 (0.006)
FNR 0.653 (0.098) 0.215 (0.121) 0.578 (0.114) 0.219 (0.116) 0.136 (0.099) 0.127 (0.097) 0.160 (0.103) 0.127 (0.096)
Accuracy 0.937 (0.009) 0.973 (0.008) 0.943 (0.010) 0.971 (0.008) 0.980 (0.007) 0.981 (0.007) 0.978 (0.007) 0.981 (0.007)
F1 1.283 (0.096) 1.694 (0.091) 1.353 (0.108) 1.679 (0.090) 1.771 (0.079) 1.780 (0.083) 1.748 (0.082) 1.785 (0.086)
MCC 0.286 (0.096) 0.702 (0.090) 0.357 (0.108) 0.688 (0.087) 0.779 (0.074) 0.788 (0.077) 0.755 (0.079) 0.792 (0.081)
$n=1000$
ERR 0.063 (0.006) 0.023 (0.006) 0.052 (0.007) 0.024 (0.006) 0.016 (0.005) 0.018 (0.005) 0.018 (0.005) 0.019 (0.005)
FPR 0.036 (0.003) 0.015 (0.004) 0.030 (0.004) 0.017 (0.004) 0.012 (0.004) 0.013 (0.004) 0.013 (0.004) 0.013 (0.004)
FNR 0.650 (0.064) 0.178 (0.084) 0.515 (0.083) 0.188 (0.085) 0.104 (0.072) 0.130 (0.072) 0.119 (0.070) 0.132 (0.074)
Accuracy 0.937 (0.006) 0.977 (0.006) 0.948 (0.007) 0.976 (0.006) 0.984 (0.005) 0.982 (0.005) 0.982 (0.005) 0.981 (0.005)
F1 1.291 (0.065) 1.744 (0.061) 1.420 (0.077) 1.727 (0.059) 1.817 (0.056) 1.792 (0.058) 1.802 (0.054) 1.788 (0.060)
MCC 0.293 (0.064) 0.749 (0.061) 0.423 (0.077) 0.732 (0.058) 0.822 (0.054) 0.796 (0.056) 0.807 (0.052) 0.793 (0.058)